### v2

Abstract
We introduce and formalize the notion of resource augmentation for maximin share allocations — an idea that can be traced back to the seminal work of Budish [JPE 2011]. Specifically, given a fair division instance with $m$ goods and $n$ agents, we ask how many copies of the goods should be added in order to guarantee that each agent receives at least their original maximin share, or an approximation thereof. We establish a tight bound of $m/e$ copies for arbitrary monotone valuations. For additive valuations, we show that at most $\min\{n-2,\left\lfloor\frac{m}{3}\right\rfloor(1+o(1))\}$ copies suffice. For approximate-MMS in instances with additive valuations, we give a tradeoff between the number of copies needed and the approximation guarantee. In particular, we prove that $\lfloor n/2\rfloor$ copies suffice to guarantee a $6/7$-approximation to the original MMS, and $\lfloor n/3\rfloor$ copies suffice for a $4/5$-approximation. Both results improve upon the best known approximation guarantees for additive valuations in the absence of copies.
1 Introduction
Fair Division goes back to antiquity, (Genesis 13), where 2 agents (Abraham and Lot) used cut and choose to split a contested resource. For divisible items and 2 agents, cut and choose ensures that (a) both agents get at least half the good, when viewed subjectively based on the agent’s own valuation — this is called a proportional allocation, and (b) no agent would prefer to swap with the other — this is called an envy free allocation. See [17, 18]. Extensions to $n>2$ agents were studied by Banach, Knaster and Steinhaus, [42], where proportionality means that every agent gets at least $1/n$ fraction of her value for the entire good. In particular, they show that proportionality is always achievable, for any number of agents, in the context of divisible cake cutting. (In general, for more than 2 agents, proportionality does not imply envy freeness.) Unfortunately, proportionality is not possible, in general, for indivisible items.
The maximin share is a relaxation of proportionality for indivisible items. Given $n$ agents and a set $M$ of indivisible items, the maximin share of an agent is the maximum value (over all partitions of $M$ into $n$ bundles), of the minimum valued bundle in the partition.
The maximin share (hereinafter denoted MMS), was studied by [32, 21]. Budish [21] defined the term and showed that if the number of agents goes to infinity and there are many copies of every item (also going to infinity) then one can approximate the maximin share to within a small error (going to zero). This was particularly relevant in the context of registration to classes, his motivating example.
Again, unfortunately, even if restricted to additive valuations, a maximin allocation need not exist [34, 27]. Albeit, there are several special cases where achieving the maximin share is in fact achievable, see [16].
Relaxations of MMS
Given that achieving the maximin share is impossible, several relaxations have been studied. Implicitly introduced in [21], a $1$-out-of-$d$ maximin share is the maximum value each agent can guarantee while partitioning the set $M$ into $d\geq n$ bundles (instead of $n$ bundles in the original MMS definition), and receiving a minimum valued bundle.
Clearly, the larger $d$ is, the easier it is to achieve the benchmark. The question then becomes, what is the smallest value of $d\geq n$ such that there is an allocation for $n$ agents so that each agent achieves the $1$-out-of-$d$ maximin share. This is also referred to as the ordinal approximation. In [33] the authors show that $1$-out-of-$\lceil 3n/2\rceil$ MMS allocations exist for any instance with additive valuations, later improved to $1$-out-of-$4\lceil n/3\rceil$ MMS allocations in [4], the current state of the art. For all we know, even $1$-out-of-$(n+1)$ MMS allocations may exist.
Another relaxation of MMS, also called approximate-MMS, is that of $\alpha$-MMS, introduced by Procaccia and Wang [34]. An $\alpha$-MMS allocation is one where every agent gets at least an $\alpha$ fraction of the maximin share. Procaccia and Wang [34] showed that $\alpha=2/3$ is attainable for additive valuations. This was subsequently improved in a sequence of papers, culminating in [1] which achieved a $(\frac{3}{4}+\frac{3}{3836})$-MMS allocation, the best known bound thus far (Akrami and Garg).
1.1 Our Contributions
Our main contribution is to introduce relaxations to the MMS by allowing resource augmentation, (i.e., allowing extra copies of items) and to obtain both positive and negative results in different settings. This relaxation was implicitly suggested by Budish in [21], in the context of course allocation (where the items are seats in a classroom), and it may also be applicable to other settings in which adding copies of an item is possible (albeit with some restrictions), e.g., assigning private tutors to students, caregivers to elderly individuals, or distributing any form of hourly-based service. We ask:
Main question:
How many items should be duplicated so as to guarantee an allocation in which every agent’s value is at least the agent’s (original) MMS?
We consider both the total number of copies and the maximum number of copies of an item. We distinguish between distinct copies (only one extra copy of an item allowed), and when more copies of a single item are allowed. Much of this paper is focused on distinct copies — up to one extra copy of an item. Our results are the following:
General Monotone Valuations
A tight bound ($\approx m/e$) on the number of copies needed to achieve MMS with copies, for arbitrary monotone valuations. More precisely we show that
  • For all $n$, $m$, there exist $n$ monotone valuations over $m$ items such that every MMS allocation with copies requires at least $\left(\frac{n-1}{n}\right)^{n}\cdot m$ copies. See Theorem 1. The number of copies of an individual good can be as high as $n$ (in an instance where $m$ is sufficiently large).
  • Given $n$ monotone valuations $\{v_{i}\}_{i\in[n]}$ and a set $[m]$ of goods, one can find an MMS allocation with copies, while making at most $\left(\frac{n-1}{n}\right)^{n}\cdot m$ extra copies. See Proposition 1. Moreover, the number of copies of any individual good is no more than $O(\ln m/\ln\ln m)$. See Proposition 2. Note that this does not conflict with the lower bound of $n$ above as the lower bound instance has $m=n^{n}$.
  • We also show a connection between two different MMS relaxations, in particular we show that if there exists a 1-out-of-$(1+\alpha)n$ MMS allocation then there is a full MMS allocation with $\lfloor\alpha m\rfloor(1+o(1))$ distinct copies, see Lemma 5. While this reduction is for general valuations — we use it to obtain bounds for additive valuations.
Additive Valuations
  • We show that
    $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$
    distinct copies suffice to guarantee a full MMS allocation, see Theorem 3. These results are indeed tight for 3 additive agents (namely, a single copy of a single good suffice), but are not known to be tight for 4 or more agents, raising a very perplexing open problem. As far as we know today, a single copy of a single good may suffice to achieve a full MMS allocation for any instance with additive valuations.
  • We also consider MMS problem for chores, where the question becomes how many chores need be removed to achieve an MMS allocation? We show that there is an MMS allocation for chores where at least $m-n+2$ of the chores will be allocated to agents, (i.e., $n-2$ chores removed), see Appendix A.
$\alpha$-MMS Approximations with Copies for Additive Valuations.
We give new approximation guarantees for additive instances, which improve upon the best-known approximations without copies. We give points along a tradeoff between the guarantee of the $\alpha$-MMS approximation and the number of copies needed. We show how to achieve a $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies (Theorem 4), and how to achieve a $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies (Theorem 5).
Relation to 1-out-of-$d$ MMS.
As mentioned, we show 1-out-of-$(1+\alpha)n$ MMS implies that exact MMS can be achieved with $\alpha m(1+o(1))$ distinct copies (see Lemma 5). In the other direction, showing that $n/\beta$ copies are sufficient for achieving exact MMS, immediately implies that 1-out-of-$(1-\beta)n$ MMS allocations exist. This is because at least $(1-\beta)n$ agents do not receive any copies, and obtain their MMS value. In particular, for additive valuations, achieving MMS with less than $n/4$ copies improves upon the currently best-known bound for 1-out-of-$d$.
1.2 Our Techniques
Our upper bounds for general monotone valuations use the probabilistic method. This is used for bounding the total number of copies as well as for bounding the maximum number of copies of a single item.
We give new extended variants of existing techniques in the fair division literature to a setting with copies. Specifically:
  • We introduce bag filling with copies. Our bag filling techniques give full MMS allocations while duplicating some of the items. See Algorithm 2: $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$.
  • Valid reductions in general take an instance and reduce it to a new instance with less agents and items, while ensuring that the MMS of each remaining agent has not decreased. An $\alpha$-valid reduction is a valid reduction where the agents removed are allocated sets whose value is at least $\alpha$ times their MMS share, see Definition 6. We give new valid reductions and $\alpha$-valid reductions that use copies, see Section 5.2.
  • While it was shown by [16] that considering agents with the same order of valuations over the items is without loss of generality when studying MMS, their proof does not carry over to the setting with copies. Nevertheless, using an augmenting path argument, we prove that for a restricted class of allocations, in particular those produced by our algorithms, such a reduction applies (see Section 5). Consequently, our algorithms for achieving approximate MMS assume, without loss of generality, that the instance is ordered. It is an intriguing open problem whether such a reduction applies in general.
  • We combine bag filling and round robin approaches to achieve approximate MMS allocations with copies for ordered instances. See Algorithm 4: $\mathsf{BagFill\mbox{-}RoundRobin}$.
1.3 Open Problems
We believe that one of our major contributions is to suggest many open problems, the penultimate open problem being “does one extra copy of a good suffice to obtain an MMS allocation for additive valuations?”. If so, this would be an interesting analogy to the Bulow Klemperer result [22] where adding a single bidder suffices so that a simple auction (Vickrey) obtains as much revenue as the optimal auction (Myerson). Other open problems are:
  • Of course, making progress on any of our bounds on additive valuations are obvious open problems, even if not showing that one extra copy suffices for MMS.
  • Considering more general classes of valuations such as submodular valuations, gross substitutes, etc. An illustrative example is that of $k$-demand valuations where we show that $n-1$ distinct copies suffice to achieve an MMS allocation, see Appendix D. We remark that the tools used to produce an MMS allocation with copies for $k$-demand valuations are a slight variant of those used for additive valuations.
  • We reduce the problem of MMS with [distinct] copies to the $1$-out-of-$d$ problem (Lemma 5). We use this to derive bounds on the number of copies in the context of additive valuations, but any such $1$-out-of-$d$ result for arbitrary monotone valuations would similarly imply analogous results for general valuations. Are there such $1$-out-of-$d$ results?
  • Another natural direction is applying the resource augmentation paradigm to other fairness notions.
1.4 Other Related Work
Fair division of indivisible goods is a vast area of research in recent years, a survey paper by Amanatidis et al., “Fair Division of Indivisible Goods: Recent Progress and Open Questions”, [7], includes 12 pages of references, and the authors stress that this is not an exhaustive survey and refer to yet other surveys (e.g., [36] with a more economic perspective) that focus on specific aspects of the problem.
Fair division notions can be roughly partitioned into share based [9] and envy based [17]. In share based fair division one ignores what the others get, one is happy if one gets value above some threshold (that does not depend on others, only on their cardinality), proportionality and MMS are share based. In envy based fair division one decides if one is happy or not depending on the others’ allocated bundles. The basic envy based fairness notion is that being envy free (one does not seek to exchange allocations with the other). For indivisible goods, envy freeness is impossible in general and thus relaxations were proposed. These include envy freeness up to one good (EF1) [35, 21] and envy freeness up to any good (EFX) [23]. EF1 allocations are known to exist, the question of when do EFX allocations exist is a notoriously hard open problem. The number of citations dealing with envy free, EF1 and EFX allocations is quite daunting, see [7]. In this paper we deal with share based, and more specifically MMS fairness.
Besides MMS and $\alpha$-MMS, other relaxations/approximations to proportionality include proportionality up to one good (PROP1), Round Robin Share, Pessimistic proportional share [24], proportional up to any good (PROPx) [8] which need not exist, and proportionality up to the maximin item (PROPm) [13] which are known to exist for 5 agents. Allocations that are both PROP1 and Pareto efficient exist [24].
Prior to Budish [21], Hill also considered the maximin share [32] and gave lower bounds on the maximin share as a function of the item values.
There was a long series of papers improving upon results for $\alpha$-MMS, complexity improvements and improvements in the quality of the result [37, 5, 34, 29, 14], $\alpha=3/4$ was achieved in [30, 28, 3] and finally [1] improved upon $3/4$ by a small constant.
A formal definition of shares and of feasible shares appears in [9, 11], MMS itself is not feasible as it may not exist, even for additive valuations. A recent share-based notion is quantile-share [12], which is guaranteed to exist for additive valuations and, more generally, for monotone valuations (then called universally feasible), assuming the Erdös Matching Conjecture holds.
The work of [2] studies EF1 and Pareto optimal allocations of chores under chore duplication, whereas our work focuses on chore deletion. In a similar spirit to our resource augmentation approach, [41] and [15] investigate the computational complexity of achieving envy-free and proportional allocations when a limited number of goods can be divided among multiple agents.
There is a long history of considering resource augmentation in various forms in computer science and economics, see chapter on resource augmentation in “Beyond the Worst-Case Analysis of Algorithms” [39]. Resource augmentation appears in Bulow Klemperer [22] where simple auctions with one extra bidder guarantee revenue at least as large as that of complex auctions without the extra bidder. This was further studied in [31] and since then has been used in many contexts such as auction design, for both revenue and welfare, prophet inequalities, etc. It also goes by the name of competition complexity, see [26, 40, 10, 19, 20, 25].
2 Model and Preliminaries
Following standard usage, define $[k]:=\{1,2,\ldots,k\}$ for integer $k\geq 1$.
An instance of a fair division problem is given by a tuple $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, where $M$ is a set of $m$ indivisible goods, and $N$ is a set of $n$ agents. Every agent $i$ has a valuation function $v_{i}:2^{M}\to\mathbb{R}_{\geq 0}$, associating a non-negative value to every set $S\subseteq M$ of goods, denoted by $v_{i}(S)$. As standard, we assume that $v_{i}$ is monotone, i.e., for any $T\subseteq S\subseteq M$, $v_{i}(T)\leq v_{i}(S)$, and normalized, so that $v_{i}(\emptyset)=0$.
A valuation function $v_{i}$ is said to be additive if for every set of items $S\subseteq M$, $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. We abuse the notation and write $v_{i}(g)$ or $v_{ig}$ instead of $v_{i}(\{g\})$ for $g\in M$.
An allocation of the set $M$ is a tuple $(A_{1},\dots,A_{n})$, where $A_{i}$ denotes the bundle allocated to agent $i$, such that $\bigcup_{i\in[n]}A_{i}=M$ (i.e., all items are allocated), and $A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$ (i.e., no item is allocated to more than one agent). We denote the collection of all allocations of the set $M$ among $n$ agents by $\mathcal{A}_{n}(M)$.
The maximin share (MMS) of agent $i$ [21] is the value that agent $i$ can guarantee by splitting the set $M$ into $n$ bundles, getting the worst out of them. Formally:
Definition 1 (Maximin share (MMS)).
For an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n$ agents, the maximin share (MMS) of agent $i$ with valuation function $v_{i}$, denoted by $\mu_{i}^{n}(M)$, is given by
$\mu_{i}^{n}(M)=\max_{A\in\mathcal{A}_{n}(M)}\min_{S\in A}v_{i}(S).$
When convenient, we sometimes abuse notation, and refer to this value as $\mu_{i}(\mathcal{I})$.
Without loss of generality, we remove all agents whose MMS value is 00. These agents get their full MMS with empty bundles. As valuations can be scaled, we may assume, without loss of generality that the MMS value for every agent is one. For additive valuations, this implies that every agent values the grand bundle (all items) at $n$ or more.
Any partition that realizes the MMS value of agent $i$, is said to be an MMS partition of agent $i$. That is, a partition $P=(P_{1},\ldots,P_{n})$ of the goods into $n$ bundles is an MMS partition of agent $i$, if and only if $\mu_{i}^{n}(M)=\min_{j}v_{i}(P_{j})$.
We next present the notion of an approximate MMS allocation.
Definition 2 ($\alpha$-MMS).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha>0$, an allocation $A$ is $\alpha$-MMS if and only if for all agents $i\in N$, $v_{i}(A_{i})\geq\alpha\cdot\mu_{i}^{n}(M)$.
Valid reductions.
Valid reductions allow us to remove a subset of agents and items while ensuring that the MMS values of the remaining agents do not decrease. The following lemma is due to [6, 16].
Lemma 1 ([6, 16]).
For any agent $i\in N$, any monotone, non-negative, valuation functions $v_{i}:2^{M}\to\mathbb{R}_{\geq 0}$ and any good $g\in M$, $\mu_{i}^{n}(M)\leq\mu_{i}^{n-1}(M\setminus\{g\}).$
By applying Lemma 1 repeatedly, we conclude that for any $k<\min\{m,n\}$, the MMS of agent $i$ may only increase if we remove an arbitrary set of $k$ goods and an arbitrary set of $k$ agents. This is shown in the following corollary.
Corollary 1.
For any agent $i\in N$, any $k<\min\{m,n\}$, and any $S\subseteq M$ s.t. $|S|=k$, it holds that
$\mu_{i}^{n}(M)\leq\mu_{i}^{n-k}(M\setminus S).$
We remark that a similar notion of $\alpha$-valid reductions is useful when pursuing approximate MMS allocations. This is defined and used in Section 5.
MMS with Copies.
An allocation with copies is an allocation where goods may be duplicated, with different instances of a same good allocated to different agents. An agent may receive at most one instance of each good. An allocation with copies is a relaxation of an allocation where the two or more bundles may have a non-empty intersection.
Definition 3 ($(t,k)$-allocation).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, a $(t,k)$-allocation $A=(A_{1},\dots,A_{n})$ is a collection of sets of items, such that
  1. 1.
    $A_{i}\subseteq M$ for every $i\in N$.
  2. 2.
    The total number of extra copies is at most $t$, i.e., $\sum_{i\in N}|A_{i}|\leq|M|+t$.
  3. 3.
    The number of extra copies of any good $g\in M$ is at most $k$, i.e., for every $g\in M$, |$|\{i\in N\mid g\in A_{i}\}|\leq k+1$.
Note that, the definition above implies that no agent is allowed to receive more than a single instance of each good.
We also consider valid reductions with copies which appear in Section 5.2.
3 MMS with Copies: General Valuations
In this section we provide (tight) bounds on the number of copies needed for full MMS under general valuations. Section 3.1 establishes bounds on the total number of copies, while Section 3.2 gives bounds on the number of copies per good.
3.1 A Tight Bound of $m/e$ on the Total Number of Copies
In this section we give a tight bound on the total number of copies required to guarantee original MMS value to all agents under monotone valuations.
Our lower bound is based on a combinatorial construction which considers $m=n^{n}$ goods positioned on the $n$-dimensional cube, so each good may be identified with $n$ coordinates, $(d_{1},\dots,d_{n})$. The valuations are such that each of the agent is almost “single-minded”, with agent $i$ only interested in bundles that contain all $n^{n-1}$ goods that share the same value in the $d_{i}$ coordinate. Those $n$ different bundles (one for each possible value of $d_{i}$), make up her MMS partition. A simple counting argument shows that the only way to satisfy all agents simultaneously is to copy $(n-1)^{n}$ goods, which is a $1/e$-fraction of the total number of goods, as $n$ tends to infinity.
We complement this result by showing, using the probabilistic method, that in every instance there exists an allocation which satisfies all agents, and copies at most $\left(1-\frac{1}{n}\right)^{n}$-fraction of the goods.
Theorem 1.
For any $n$, there exists an instance with $n$ monotone valuations such that the total number of copies needed is $\left(\frac{n-1}{n}\right)^{n}$-fraction of the goods, and there is at least one good which must be copied $n-1$ times.
Proof.
Consider $n$ agents and $n^{n}$ goods positioned on the $n$-dimensional cube, i.e., the set of goods is
$M=\{g_{d_{1},\dots,d_{n}}\mid d_{1},\dots,d_{n}\in[n]\}.$
Each agent $i\in[n]$ is interested in one of $n$ disjoint bundles, $P_{i}^{1},\dots,P_{i}^{n}$ of $n^{n-1}$ goods each, where $P_{i}^{j}=\{g_{d_{1},\dots,d_{n}}\mid d_{i}=j\}$. Formally, agent $i$’s valuation is monotone and for every $S\subseteq M$ it is defined to be:
$v_{i}(S)=\begin{cases}1&\exists j\in[n]\text{ s.t. }P_{i}^{j}\subseteq S\\ 0&\text{otherwise}\end{cases}$
Agent $i$’s MMS partition is $(P_{i}^{1},\dots,P_{i}^{n})$ and her MMS value is 1.
Observe that for any $k$ agents, $i_{1},\dots,i_{k}\in[n]$ and for any $k$ indexes $j_{1},\dots,j_{k}\in[n]$, it holds that
$|P_{i_{1}}^{j_{1}}\cap P_{i_{2}}^{j_{2}}\cap\dots\cap P_{i_{k}}^{j_{k}}|=|\{g_{d_{1},\dots,d_{n}}\mid\forall l\in[k]\;\;d_{i_{l}}=j_{l}\}|=n^{n-k}.$
If we wish to find an allocation with copies $A=(A_{1},\dots,A_{n})$ such that for every $i$, it holds that $v_{i}(A_{i})\geq 1$, while minimizing the number of items we copy, then it is without loss to assume that $A_{i}=P_{i}^{j_{i}}$ for some $j_{i}\in[n]$. Thus, the total number of goods (copies included) is $\sum_{i\in[n]}|A_{i}|=\sum_{i\in[n]}|P_{i}^{j_{i}}|=n\cdot n^{n-1}=n^{n}$. The number of different goods used is, using the inclusion-exclusion principle:
$\left|\bigcup_{i\in[n]}P_{i}^{j_{i}}\right|$ $=$ $\sum_{k=1}^{n}(-1)^{k+1}\binom{n}{k}n^{n-k}=n^{n}-n^{n}-\left(\sum_{k=1}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
$=$ $n^{n}-\left(\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)=n^{n}-(n-1)^{n},$
which implies that the total number of copies is $(n-1)^{n}$. Additionally, in any such allocation there exists a good which is allocated $n$ times, and thus duplicated $n-1$ times, as $g_{j_{1},j_{2},\dots,j_{n}}\in\bigcap_{l=1}^{n}P_{i_{l}}^{j_{l}}$, and the claim follows. ∎
Below we prove a matching upper bound on the number of total copies required. We show that there always exists an allocation which assigns to every agent $i$ one of his MMS bundles and makes no more than $\left(1-\frac{1}{n}\right)^{n}\cdot m$ additional copies.
Proposition 1.
For any set $N$ of $n$ agents with monotone valuations $\{v_{i}\}_{i\in N}$, and any set $M$ of $m$ goods, one can find an MMS allocation with copies, while making at most $\left(\frac{n-1}{n}\right)^{n}\cdot m$ extra copies.
Proof.
Fix $n$ MMS partitions $\{(P_{i}^{1},\dots,P_{i}^{n})\}_{i\in N}$. Let $A=(A_{1},\dots,A_{n})$ be a random allocation, where $A_{i}\sim U(\{P_{i}^{1},\dots,P_{i}^{n}\})$, independently for any agent $i$. That is, agent $i$ gets one of her MMS bundles chosen uniformly at random.
Observe that for every agent $i$ and good $g$, $\Pr[g\in A_{i}]=\frac{1}{n}$, and so the expected number of unique items in allocation $A$ is:
$\mathbb{E}_{A}\left[|\cup_{i\in[n]}A_{i}|\right]$ $=$ $\sum_{g\in[m]}\Pr[g\in\cup_{i\in[n]}A_{i}]=\sum_{g\in[m]}(1-\Pr[g\notin\cup_{i\in[n]}A_{i}])=\sum_{g\in[m]}\left(1-\left(1-\frac{1}{n}\right)^{n}\right)$
$=$ $m\cdot\left(1-\left(1-\frac{1}{n}\right)^{n}\right),$
where the third inequality follows from independence.
The expected total number of items (including copies) is
$\mathbb{E}_{A}\left[\sum_{i=1}^{n}|A_{i}|\right]=\sum_{i=1}^{n}\mathbb{E}_{A}\left[|A_{i}|\right]=n\cdot\frac{m}{n}=m.$
By linearity of expectation, the expected number of copies made for an allocation $A$ is
$\mathbb{E}_{A}\left[\sum_{i=1}^{n}|A_{i}|-|\cup_{i\in[n]}A_{i}|\right]=m-m\cdot\left(1-\left(1-\frac{1}{n}\right)^{n}\right)=m\left(1-\frac{1}{n}\right)^{n},$
so there must exists an allocation that assigns each agent one of her MMS bundles and requires no more than $m\cdot\left(1-\frac{1}{n}\right)^{n}$ extra copies. ∎
3.2 Bound on the Maximum Number of Copies of Any Good
In this section we upper bound the number of copies of any individual good, and together with the bound established in Section 3.1, show to achieve both guarantees with high probability.
Theorem 2.
For any instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with monotone valuations, there exists a $(t,k)$-allocation that satisfies all agents and satisfies $t\leq m/e$ and $k\leq\min\left\{n-1,\frac{3\ln m}{\ln\ln m}\right\}$. Moreover, given an MMS partition for each agent, for any $\beta\in\mathbb{N}$, there exists an efficient algorithm that returns such an allocation with probability $1-e^{-\beta}$.
In the proof of Theorem 2, we use the following lemma. Its proof, which relies on a balls-and-bins argument, is deferred to Appendix B.
Lemma 2.
For any fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, with monotone valuations $\{v_{i}\}_{i\in N}$, there exists an MMS allocation where every good is copied at most $O\left(\frac{\ln m}{\ln\ln m}\right)$ times.
Proof sketch.
Consider an allocation where agent $i$’s bundle, is randomly chosen as one of the sets in $i$’s MMS partition. We view this process as throwing $n$ balls (agents) into $n$ bins (the index in an MMS partition). Fix some good $g\in M$, it belongs to some MMS set for all the agents, possibly different indices, but we can rename the sets and assume without loss of generality that $g$ is always in the set of index 1 in the MMS partitions of all agents. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. For $k=\frac{3\ln m}{\ln\ln m}$, the probability of having a bin with more than $k$ balls is at most $\frac{1}{m}$. ∎
Proof of Theorem 2.
Fix an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$. For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be her MMS partition. Consider the algorithm $\mathcal{A}$ which does the following (at most) $3m\beta$ times:
  1. 1.
    Samples an allocation according to $A_{i}\sim U(P^{i})$.
  2. 2.
    Returns the allocation $A$ only if the number of total copies is $t\leq m/e$ and every item is copied at most $k\leq\frac{3\ln m}{\ln\ln m}$.
First, by the union bound, the probability that $\mathcal{A}$ fails in some iteration is
[…]
where the last inequality follows from the proof of Lemma 2.
Recall that in the proof of Proposition 1, we have established that when the allocation $A$ is picked uniformly at random, as in the above procedure, the expected number of total copies is $m/e$. This allows us to bound the second term using Markov’s inequality,
$\Pr[t>m/e]\leq\Pr\left[t\geq\left(1+\frac{e}{m}\right)\frac{m}{e}\right]\leq\frac{1}{1+\frac{e}{m}}\leq\frac{m}{m+2}.$
Putting everything together we have,
[…]
In particular, there is a strictly positive probability that $\mathcal{A}$ succeeds in some iteration, so there exists an MMS allocation with $t\leq m/e$ and $k\leq\frac{3\ln m}{\ln\ln m}$. The statement follows immediately, as
[…]
4 MMS with Copies: Additive Valuations
In this section, we prove the following theorem.
Theorem 3.
For any instance with additive valuations, at most $\min\{n-2,\lfloor m/3\rfloor(1+o(1))\}$ distinct copies suffice to guarantee a full MMS allocation.
The proof of Theorem 3 is divided into two parts. In Section 4.1, we give an algorithm that obtains full MMS with $n-2$ distinct copies. In Section 4.2 we relate the $1$-out-of-$d$ bounds to our problem (a relation that holds for arbitrary monotone valuations), and use it to obtain full MMS with $\lfloor m/3\rfloor(1+o(1))$ distinct copies for additive valuations.
4.1 Full MMS with $n-2$ Copies
Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ operates as follows. It inspects the MMS Partition $P$ of an arbitrary agent, tries to satisfy as many agents as possible (without harming other agents) using $P$ without any copies. Then it uses a modified bag-filling algorithm, $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$, that uses copies in order to satisfy the remaining agents. Recall that we scale valuations so that $\mu_{i}^{n}(M)=1$ for every agent $i$, which also implies $v_{i}(M)\geq n$.
while $\exists i\in N,j\in M$ s.t. $v_{ij}\geq 1$ do
       $A_{i}\leftarrow\{j\}$$N\leftarrow N\setminus\{i\}$, $M\leftarrow M\setminus\{j\}$
end while
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be the MMS partition of agent 1
Consider the bipartite graph $G=(V=N\times P,E)$, where $E=\{(i,P_{j})\ :\ v_{i}(P_{j})\geq 1\}$
if there is a perfect matching between $N$ and $P$ in $G$ then
       Allocate each $i\in N$ its match in the perfect matching
end if
else
       Let $P^{\prime}\subset P$ be the minimal set of bundles violating Hall’s theorem ($|P^{\prime}|>1$ since $1$ is matched to every node in $P$)
       Consider some $P_{j}\in P^{\prime}$ and let $\hat{P}=P^{\prime}\setminus\{P_{j}\}$
       Find a matching between bundles in $\hat{P}$ and agents in $N$
       foreach $i\in N$ matched to some $P_{j}\in\hat{P}$ do
             $A_{i}\leftarrow P_{j}$
       end foreach
      Let $\tilde{N}$ be the set of unallocated agents and $\tilde{M}=M\setminus\left(\bigcup_{i\in N\setminus\tilde{N}}A_{i}\right)$ the set of unallocated items
       Allocate agents in $\tilde{N}$ using procedure $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}((\tilde{N},\tilde{M},\{v_{i}\}_{i\in\tilde{N}}))$
end if
Input: $n$ additive valuations $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-2$ distinct copies.
Algorithm 1 $\mathsf{Match\mbox{-}n\mbox{-}Fill}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive valuations $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-2$ distinct copies.
$t\leftarrow 0$, $B_{t}\leftarrow\emptyset$, $N_{0}\leftarrow N$, $M_{0}\leftarrow M$
for items $j\leftarrow 1$ to $m$ do
       if $\exists i_{t}\in N_{t}$ such that $v_{i_{t}}(B_{t}\cup\{j\})\geq 1$ then
             $A_{i_{t}}\leftarrow B_{t}\cup\{j\}$
             $t\leftarrow t+1$, $B_{t}\leftarrow\{j\}$
             $N_{t}\leftarrow N_{t-1}\setminus\{i_{t-1}\}$, $M_{t}\leftarrow M_{t-1}\setminus B_{t-1}$
       end if
      else
             $B_{t}\leftarrow B_{t}\cup\{j\}$
       end if
end for
Input: $n$ additive valuations $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-1$ distinct copies.
Algorithm 2 $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive valuations $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at most $n-1$ distinct copies.
We first show that $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ satisfies all agents, assuming each agent values the remaining items high enough (proof in Appendix C). An example of a run of $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ appears in Appendix E.
Lemma 3.
Assuming that for every $i\in N$, $v_{i}(M)\geq|N|$ and that for every item $j$, $v_{ij}<1$, $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ outputs an allocation such that $v_{i}(A_{i})\geq 1$ for every $i\in N$, and creates at most $|N|-1$ distinct copies.
Proof sketch..
A simple induction shows that every time we allocate a bundle, the value of every remaining agent $i$ for all remaining items is at least the number of remaining agents times $i$’s MMS value. Indeed, $i$’s value for all items except the last item inserted to the bag is at most $1$, otherwise, $i$ could have been allocated in a previous iteration. Since we duplicate the last inserted item, we get that the number of agents decreased by 1, while the total value of $i$ for all remaining items decreased by at most 1. Thus, we make sure that at any point, we have enough value to satisfy all remaining agents. ∎
We now show that $n-2$ distinct copies suffice.
Lemma 4.
For any fair division instance with additive valuations, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ guarantees a full MMS allocation for all the agents with at most $n-2$ distinct copies.
Proof.
First, by definition, every agent that is assigned some bundle for which they have an edge in $G$ gets their MMS value. Moreover, since agent $1$ has an edge to all bundles in $P$, then at least one agent gets satisfied, and $|\tilde{N}|\leq|N|-1$. It remains to show that sets $\tilde{N}$ and $\tilde{M}$ satisfy the conditions of Lemma 3.
$\mathsf{Match\mbox{-}n\mbox{-}Fill}$ first allocates items with value more than the MMS of an agent, so for every $i\in\tilde{N},j\in\tilde{M}$, $v_{ij}<1$. Moreover, by Lemma 1, allocating a single item to a single agent is a valid reduction. Thus, after this step, we have that for every $i\in N$, $v_{i}(M)\geq|N|$. We show that after allocating bundles $\hat{P}$ to agents in $N$, for every unallocated agent $i\in\tilde{N}$, $v_{i}(\tilde{M})\geq|\tilde{N}|$.
Recall that $P^{\prime}$ is a minimal set of bundles violating Hall’s Theorem, and let $|P^{\prime}|=k$. Let $N^{\prime}=\Gamma(P^{\prime})$ be their neighbors in $G$. Note that $|N^{\prime}|=k-1$, otherwise, one can remove an arbitrary element of $P^{\prime}$ and get a smaller set violating Hall’s Theorem. Moreover, since $\hat{P}$ does not violate Hall’s theorem, we know that $\Gamma(\hat{P})=N^{\prime}$ and that the agents in $N^{\prime}$ are exactly the agents matched by bundles from $\hat{P}$. Thus, it must be that there are no edges between agents in $\tilde{N}=N\setminus N^{\prime}$ and bundles in $\hat{P}$. Therefore, for every $i\in\tilde{N}$, $v_{i}(\bigcup_{S\in\hat{P}}S)\leq k-1$, which implies
$v_{i}(\tilde{M})=v_{i}(M)-v_{i}\left(\bigcup_{S\in\hat{P}}S\right)\geq|N|-(k-1)=|N|-|N^{\prime}|=|\tilde{N}|,$
as desired. It follows that all the conditions required in order to apply Lemma 3 are fulfilled, and we can satisfy agents in $\tilde{N}$ using $|\tilde{N}|-1\leq|N|-2\leq n-2$ distinct copies (recall that at least one agent was already allocated a bundle before). Since $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is the only place where we produce copies, this concludes the proof. ∎
Together with the example in [27] of an instance with $n=3$ additive agents that does not admit an MMS allocation, Lemma 4 implies a tight bound on the number of copies for $3$ additive agents.
Corollary 2.
For every instance with 3 additive valuations, a single copy of a single item suffices to guarantee a full MMS allocation. This is tight, i.e., there are instances where one copy is required.
Remark.
One can consider a similar question in the context of chore allocation, and ask “what is the maximum number of chores that can be assigned to agents such that no agent’s cost is larger than their maximin share”; that is, how many chores are needed to be set aside in order to produce a full MMS allocation. It turns out that an analogous process to $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ in the context of chores outputs an allocation where at most $n-2$ items are disposed of, and each agent is allocated a bundle of cost no larger than their MMS value. More on this is in Appendix A.
4.2 Full MMS with $\lfloor m/3\rfloor(1+o(1))$ Copies
We now get an improved bound when $m$ is not too large. This is done via a reduction from 1-out-of-$(1+\alpha)n$ bounds.
Lemma 5.
If there exists a 1-out-of-$(1+\alpha)n$ MMS allocation for every fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ for some constant $\alpha\in(0,1)$ then
$\lfloor\alpha m\rfloor+\left\lceil\frac{(1+\alpha)^{2}}{n-1-\alpha}m\right\rceil=\lfloor\alpha m\rfloor(1+o(1))$
distinct copies suffice to guarantee a full MMS allocation. This holds for arbitrary monotone valuations.
Proof.
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance for which there exists a 1-out-of-$(1+\alpha)n$ MMS. Notice that if we take a set of agents $N^{\prime}$ of size $n^{\prime}$ for which $(1+\alpha)n^{\prime}\leq n$ then we can use the guarantee stated in the lemma to produce an allocation $\{A_{i}\}_{i\in N^{\prime}}$ without copies such that for each agent $i\in N^{\prime}$,
$v_{i}(A_{i})\geq\mu_{i}^{\lceil(1+\alpha)n^{\prime}\rceil}(M)\geq\mu_{i}^{n}(M).$
Consider an arbitrary set $N^{\prime}$ of $n^{\prime}=\lfloor n/(1+\alpha)\rfloor$ agents and the allocation that satisfies their MMS value, $\{A_{i}\}_{i\in N^{\prime}}$. Consider the set $\tilde{N}=N\setminus N^{\prime}$. We proceed as follows.
Consider a set
$S^{\prime}\in\operatorname*{arg\,min}_{S\subseteq N^{\prime}\ :\ |S|=|\tilde{N}|}|\cup_{i\in S}A_{i}|.$
Allocate each agent $i\in S^{\prime}$ the set $A_{i}$. Now consider the agents in $N\setminus S^{\prime}$. This is a set of size $\lceil n/(1+\alpha)\rceil,$ and therefore there exists an allocation $\{B_{i}\}_{i\in N\setminus S^{\prime}}$ of the set $M$ of items without copies which guarantees each agent in $N\setminus S^{\prime}$ their MMS value. Allocate each agent $i\in N\setminus S^{\prime}$ the set $B_{i}$. Notice that this allocation satisfies the MMS of all agents, and the only copies made are the items in $\cup_{i\in S^{\prime}}A_{i}$. We next bound the number of items in this set.
First, notice that
$|S^{\prime}|=|\tilde{N}|=n-\lfloor n/(1+\alpha)\rfloor=\left\lceil\frac{\alpha}{1+\alpha}n\right\rceil.$
Since we chose the set of $|\tilde{N}|$ agents with the least number of items in $N^{\prime}$, it holds that
$|\cup_{i\in S^{\prime}}A_{i}|$$\leq$ $\left\lfloor\frac{|\tilde{N}|}{|N^{\prime}|}m\right\rfloor=\left\lfloor\frac{\lceil\frac{\alpha}{1+\alpha}n\rceil}{\left\lfloor\frac{1}{1+\alpha}n\right\rfloor}m\right\rfloor\leq\left\lfloor\frac{\frac{\alpha}{1+\alpha}n+1}{\frac{1}{1+\alpha}n-1}m\right\rfloor$
$=$ $\left\lfloor\left(\alpha+\frac{1+\alpha}{\frac{1}{1+\alpha}n-1}\right)m\right\rfloor$
$\leq$ $\lfloor\alpha m\rfloor+\left\lceil\frac{(1+\alpha)^{2}}{n-1-\alpha}m\right\rceil$
$=$ $\lfloor\alpha m\rfloor(1+O(1/n))=\lfloor\alpha m\rfloor(1+o(1)).$
We apply the above lemma to the following state-of-the-art bound due to [4].
Lemma 6 (Akrami et al. [4]).
For every fair division instance with additive valuations, there exist a 1-out-of-$4\lceil n/3\rceil$ MMS allocation.
Thus, we get the following.
Corollary 3.
For any fair division instance with additive valuations, at most $\lfloor\frac{m}{3}\rfloor(1+o(1))$ distinct copies suffice to guarantee a full MMS allocation.
Proof.
Since $4\lceil n/3\rceil\leq 4n/3+4=n(1+\frac{1}{3}+\frac{4}{n})$, applying Lemma 5 with $\alpha=\frac{1}{3}+\frac{4}{n}$ implies that the number of copies sufficient in order to produce a full MMS allocation is at most
$\left\lfloor\left(\frac{1}{3}+\frac{4}{n}\right)m\right\rfloor+\left\lceil\frac{(1+\frac{1}{3}+\frac{4}{n})^{2}}{n-1-\frac{1}{3}-\frac{4}{n}}m\right\rceil$$\leq$ $\left\lfloor\frac{m}{3}\right\rfloor+\left\lceil\frac{4}{n}m\right\rceil+\left\lceil\frac{(1+\frac{1}{3}+\frac{4}{n})^{2}}{n-\frac{4}{3}-\frac{4}{n}}m\right\rceil$
$=$ $\left\lfloor\frac{m}{3}\right\rfloor(1+O(1/n))$
$=$ $\left\lfloor\frac{m}{3}\right\rfloor(1+o(1)).$
5 Approximate-MMS with Copies: Additive Valuations
In this section we study approximate MMS allocations while allowing duplication of goods.
First, in Section 5.1, we convert a fair division instance into an ordered instance with $n$ agents and $m$ items, in which all agents have the same ordinal preferences.
It has been observed, for the original MMS problem (without copies), that it is without loss of generality to consider ordered instances [16, 14]. Notably, where copies are allowed, the same transformation as in [16, 14] may no longer yield a valid allocation, see example in Appendix F. To address this, we impose an additional constraint on the allocation in the ordered instance (see Definition 5), and show that under this constraint, it suffices to consider ordered instances. We then devise algorithms that produce allocations adhering to this constraint.
In Section 5.2, we review some of the known reduction rules and introduce new ones. Then, in Section 5.3, we present a simple algorithm $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies (see Algorithm 4). This algorithm is the building block for achieving $6/7$-MMS and $4/5$-MMS with copies in the sections that follow. In particular, in Sections 5.4 and 5.5, we obtain $6/7$-MMS with $\lfloor n/2\rfloor$ distinct copies and $4/5$-MMS with $\lfloor n/3\rfloor$ distinct copies, respectively, by first reducing the initial instance using appropriate valid reductions, and then applying $\mathsf{BagFill\mbox{-}RoundRobin}$.
5.1 Ordered Instance
Given an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, for every agent $i\in N$, let $\sigma_{i}:[m]\to[m]$ be a permutation that orders the goods according to $i$’s preferences, namely $\sigma_{i}(j)$ is the $j$th most valued good (breaking ties by index111any consistent tie breaking rule will do., and thus $v_{i}(\sigma_{i}(1))\geq v_{i}(\sigma_{i}(2))\geq\dots\geq v_{i}(\sigma_{i}(m))$.
As done by Bouveret and Lemaître [16], we construct an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ with a new set of “virtual goods”, $M^{\prime}=\{1,\dots,m\}$, and modified valuations, $\{v^{\prime}_{i}\}_{i\in N}$, such that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$ for any $j\in M^{\prime}$.
Definition 4 (Ordered Counterpart ([16])).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, its ordered counterpart, $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, is an instance with the same set of agents and a set of “virtual goods”, $M^{\prime}$, of the same cardinality as $M$. For each agent $i\in[n]$, and any $j\in M^{\prime}$, it holds that $v^{\prime}_{i}(j)=v_{i}(\sigma_{i}(j))$.
Observe that by definition, all agent in $\mathcal{I}^{\prime}$ have the same ordinal preferences, i.e., for every $i\in N$, $v^{\prime}_{i}(1)\geq v^{\prime}_{i}(2)\geq\dots\geq v^{\prime}_{i}(m)$. It was shown in [16] that to guarantee agent $i$ a value of at least $\alpha_{i}$ in an instance $\mathcal{I}$, it suffices to find an allocation providing this guarantee in its ordered counterpart $\mathcal{I}^{\prime}$. This is cast in the following proposition.
Proposition 2 (Reduction to ordered without copies [16]).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Any allocation (without copies) $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ for $\mathcal{I}^{\prime}$ can be converted in poly-time to an allocation (without copies) $A=(A_{1},\dots,A_{n})$ for $\mathcal{I}$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
Proof.
The transformation from the ordered allocation $A^{\prime}$ to the allocation $A$ is done via a simple picking sequence procedure: In every iteration $j\in\{1,\dots,m\}=M^{\prime}$, let $i$ be the agent such that $j\in A^{\prime}_{i}$. Agent $i$ chooses the most-valuable item in $M$ (according to $v_{i}$), which has not yet been chosen.
Since by iteration $j$ exactly $j-1$ items have been chosen, agent $i$ always chooses a good $g\in M$ such that $v_{i}(g)\geq v_{i}(\sigma_{i}(j))=v^{\prime}_{i}(j)$ and thus, by additivity, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. ∎
We will show a similar result to Proposition 2, for allocations with $t$ distinct copies in which the top $t$ goods (and their copies) are allocated to $2t$ different agents.
Definition 5 (Simple Allocation).
Given an ordered instance $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$, an allocation with $t$ total copies $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ is called simple if
  • There is at most one copy of each good.
  • The duplicated items are $g_{1},\dots,g_{t}$ for $t\leq n/2$.
  • There exists a set of $2t$ distinct agents, $i_{1},\dots,i_{t},i^{\dagger}_{1},\dots,i^{\dagger}_{t}$ such that for any $j\in[t]$, agents $i_{j}$ and $i^{\dagger}_{j}$ are allocated the two instances of $g_{j}$, i.e., $A^{\prime}_{i_{j}}\cap A^{\prime}_{i^{\dagger}_{j}}=\{g_{j}\}$.
Recall that agent $i$ cannot get two instances of the same good in $A^{\prime}_{i}$. [^2] Otherwise, and if additive valuations are extended so that two instances of the same good yield twice its value, then the same argument as in Proposition 2 would have implied the correctness of Proposition 3.
Proposition 3 (Reduction to ordered for simple allocations).
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance and let $\mathcal{I}^{\prime}=(N,M^{\prime},\{v^{\prime}_{i}\}_{i\in N})$ be its ordered counterpart. Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation for $\mathcal{I}^{\prime}$ with $t$ distinct copies. Then, $A^{\prime}$ can be converted in poly-time to an allocation for $\mathcal{I}$ with $t$ distinct copies, $A=(A_{1},\dots,A_{n})$, such that for every agent $i$, $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$.
In order to prove Proposition 3, we presented the following useful lemma.
Lemma 7.
Fix an agent $i\in N$ and a threshold $\tau\in\mathbb{R}_{+}$. Let $X,Y\subset M$ be such that
  • $|X|=|Y|=k$, $X=\{x_{1},\ldots,x_{k}\}$, $Y=\{y_{1},\ldots,y_{k}\}$, and
  • for all $j\in[k]$, $v_{i}(y_{j})\geq v_{i}(x_{j})$.
If there exists a set $G\subset M\setminus X$ of $\ell$ distinct goods such that $v_{i}(X\cup\{g\})\geq\tau$ for all $g\in G$, then there exists a set $H\subset M\setminus Y$ of $\ell$ distinct goods such that $v_{i}(Y\cup\{h\})\geq\tau$ for all $h\in H$.
Proof.
Observe that by additivity, any good $g\in G\setminus Y$ satisfies $v_{i}(Y\cup\{g\})\geq v_{i}(X\cup\{g\})\geq\tau$. Thus, it suffices to show that for every $g\in G\cap Y$, there exists a (distinct) corresponding good $h(g)\in M\setminus(Y\cup G)$ such that $v_{i}(Y\cup\{h(g)\})\geq\tau$. If this requirement is met, the set $H=(G\setminus Y)\cup\{h(g)\mid g\in G\cap Y\}$ satisfies the desired condition in the lemma. We construct the mapping from every $g$ to $h(g)$ using an augmenting path argument in a bipartite graph (see Figure 1 for an illustration).
Consider the directed bipartite graph $\mathcal{G}=(\mathcal{X},\mathcal{Y},\mathcal{E})$, where $\mathcal{X}=\{\mathcal{X}_{g}\}_{g\in X}$ and $\mathcal{Y}=\{\mathcal{Y}_{g}\}_{g\in Y}$. Every vertex $\mathcal{X}_{g}\in\mathcal{X}$ (respectively, $\mathcal{Y}_{g}\in\mathcal{Y}$ ) corresponds to good $g\in X$ (resp., $g\in Y$). Note that $|\mathcal{X}|=|\mathcal{Y}|=k$. When clear in the context, we abuse notation and write $\mathcal{X}_{j}$ instead of $\mathcal{X}_{x_{j}}$ (and similarly for $\mathcal{Y}$).
For every $j\in[k]$, we add the (black) edge $(\mathcal{Y}_{j},\mathcal{X}_{j})$. In addition, for any good $g\in X\cap Y$, we add the (red) edge $(\mathcal{X}_{g},\mathcal{Y}_{g})$. Observe that $\mathcal{E}$ is a collection of disjoint cycles of length 2, or paths333 Assume that both $X$ and $Y$ are ordered according to $i$’s valuation: $v_{i}(x_{1})\geq\dots\geq v_{i}(x_{k})$ and the same for $Y$, where tie breaking is consistent among $X$ and $Y$. Observe that path between $\mathcal{Y}_{y_{j}}$ and $\mathcal{Y}_{y_{j^{\prime}}}$ implies that agent $i$ prefers $y_{j}$ over $y_{j^{\prime}}$, so a cycle of length greater than 2 is not possible. . Additionally, every vertex $v\in\mathcal{X}\cup\mathcal{Y}$ satisfies $deg(v)\in\{1,2\}$, where by $deg$ we consider both in- and out-degrees.
Fix $g\in G\cap Y$. Since $g\in G$, then $g\notin X$ and $deg(\mathcal{Y}_{g})=1$. Let $P=(\mathcal{Y}_{g},\mathcal{X}_{g_{1}},\mathcal{Y}_{g_{1}},\dots,\mathcal{X}_{g_{\ell}},\mathcal{Y}_{g_{\ell}},\mathcal{X}_{g^{\prime}})$ be the unique path originating at $\mathcal{Y}_{g}$. We claim that $h(g)=g^{\prime}$ satisfies the desired requirement.
Since $\mathcal{X}_{g^{\prime}}$ has no outgoing edges, $g^{\prime}\notin Y$. Also, since $g^{\prime}\in X$, then $g^{\prime}\notin G$. Thus, $g^{\prime}\in M\setminus(Y\cup G)$. It remains to show that $v_{i}(Y\cup\{g^{\prime}\})\geq\tau$. To this end, let $Q=\{g_{1},\dots,g_{\ell}\}\subseteq X\cap Y$, as defined by the path $P$. Observe that the set of goods which correspond to $\mathcal{Y}\cap P$ is exactly $Q\cup\{g\}$ (see Figure 1). Similarly, the set of goods corresponding to $\mathcal{X}\cap P$ is exactly $Q\cup\{g^{\prime}\}$. Since every vertex in $\mathcal{Y}\setminus P$ is pointed to by exactly one vertex in $\mathcal{X}\setminus P$, and $v_{i}(y_{j})\geq v_{i}(x_{j})$ for every $j\in[k]$, it follows that $v_{i}(Y\setminus(\{Q\cup\{g\}))\geq v_{i}(X\setminus(\{Q\cup\{g^{\prime}\}))$. By the last inequality and additivity of $v_{i}$ we get
$v_{i}(Y\cup\{g^{\prime}\})$ $=v_{i}(Y\setminus(Q\cup\{g\}))+v_{i}(Q\cup\{g\})+v_{i}(\{g^{\prime}\})$
$\geq v_{i}(X\setminus(Q\cup\{g^{\prime}\}))+v_{i}(Q\cup\{g^{\prime}\})+v_{i}(\{g\})$
$=v_{i}(X\cup\{g\})$
$\geq\tau,$
which concludes the proof.
[…]
Figure 1: Illustration of the augmenting path argument in the proof of Lemma 7. The element $g\in G$ (whose corresponding node is $\mathcal{Y}_{g}$, $g=y_{1}$) is replaced with $g^{\prime}$ (whose corresponding node is $\mathcal{X}_{g^{\prime}}=\mathcal{X}_{4}$, $g^{\prime}=x_{4}$). We write $\mathcal{X}_{i}$ / $\mathcal{Y}_{i}$ rather than the cumbersome $\mathcal{X}_{x_{i}}$/$\mathcal{Y}_{y_{i}}$. It is purely an example that the path nodes are associated with a contiguous prefix of the goods, e.g., $\mathcal{X}_{1}$ could point to $\mathcal{Y}_{5}$.
Proof of Proposition 3.
Let $A^{\prime}=(A^{\prime}_{1},\dots,A^{\prime}_{n})$ be a simple allocation in which the goods $g_{1}<\ldots<g_{t}$ for $t\leq n/2$ are copied. Rename the agents so that for every $j\in[t]$ one copy of $g_{j}$ is allocated to agent $j$, i.e., $g_{j}\in A_{j}$. We will refer to each such $g_{j}$ as the ‘copy’, and to the other instance of $g_{j}$ (assigned to some agent $i\in N\setminus\{1,\dots,t\}$), as the ‘original’.
We will use a similar picking sequence procedure to the one given in Proposition 2, but with two rounds. First, we will only use the $m$ ‘original’ goods to determine the picking sequence, this is identical to the picking sequence presented earlier. Then, another picking sequence will use the ‘copies’ (and so will involve agents $1,\dots,t$, where each picks one good). Each agent in the second round may pick any good in $M$ which was not yet allocated in this round and was not picked by him in the first round. Let $A_{i}\subseteq M$ be the set of goods picked by agent $i$ in both rounds.
Any agent $i\in N\setminus\{1,\dots,t\}$ only participates in the first round, so as in Proposition 2, it holds that $v_{i}(A_{i})\geq v^{\prime}_{i}(A^{\prime}_{i})$. For agent $j\in\{1,\dots,t\}$, let $B_{j}$ be the set of $|A_{j}|-1$ goods he picked in the first round. To show that $v_{j}(A_{j})\geq v^{\prime}_{j}(A^{\prime}_{j})$, it is enough to show that at the beginning of the second round there is a set of goods $H\subseteq M\setminus B_{j}$, such that (i) $|H|\geq j$, (ii) each $h\in H$ satisfies $v_{j}(B_{j}\cup\{h\})\geq v^{\prime}_{j}(A^{\prime}_{j})$. This would imply that in round $j$, player $j$ can pick a good $h$ such that $A_{j}=B_{j}\cup\{h\}$ satisfies the claim.
Denote $A^{\prime}_{j}=\{j,j_{1},\dots,j_{k}\}$, $B_{j}=\{b_{1},\dots,b_{k}\}$, where $b_{\ell}$ was picked by agent $j$ in round $j_{\ell}$. The existence of $H$ follows from applying Lemma 7 with $X=\{\sigma_{j}(j_{1}),\dots,\sigma_{j}(j_{k})\}$, $Y=B_{j}$, $G=\{\sigma(g_{1}),\dots,\sigma_{j}(g_{j})\}$ and $\tau=v^{\prime}_{j}(A^{\prime}_{j})$. Observe that by definition of the first picking sequence, for any $\ell$, it holds that $v_{i}(b_{\ell})\geq v_{i}(\sigma_{j}(j_{\ell}))=v^{\prime}_{j}(j_{\ell})$. Also, $\tau=v^{\prime}_{j}(A^{\prime}_{j})=v^{\prime}_{j}(A^{\prime}_{j}\setminus\{g_{j}\})+v^{\prime}_{j}(g_{j})=v_{j}(X)+v_{j}(\sigma_{j}(g_{j}))=v_{j}(X\cup\{\sigma_{j}(g_{j})\})$. By definition, any good in $g\in G$ has $v_{j}(\{g\})\geq v_{j}(\sigma_{j}(g_{j}))$, and by additivity all the conditions of Lemma 7 are satisfied. This concludes the proof. ∎
For the rest of this section, we only consider ordered instances. Because the algorithms we use to solve these instances always output a simple allocation, and therefore, their guarantees, by Proposition 3, can be extended to any instance, even unordered.
5.2 Reduction Rules
Definition 6.
Given an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, an allocation rule that allocates bundles $A_{1},\ldots,A_{k}$ to agents $1,\ldots,k$ respectively is $\alpha$-valid, if and only if the following holds:
  • $v_{i}(A_{i})\geq\alpha\cdot\mu_{i}(\mathcal{I})$ for all $i\in[k]$, and
  • $\mu^{n-k}_{i}(M\setminus(A_{1}\cup\ldots\cup A_{k}))\geq\mu^{n}_{i}(M)$ for all $i\in N\setminus[k]$.
The following lemma with small value of $k$ (namely $k\leq 3$) has been proven and used in [28, 3, 1].
Lemma 8.
Given an ordered fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, a fixed agent $i\in N$, and integer $k<|M|/n$, let $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$. Then, $\mu^{n-1}_{i}(M\setminus K)\geq\mu^{n}_{i}(M).$
Proof.
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be an MMS partition of agent $i$ for the original instance. By the pigeonhole principle, there exists $j\in[n]$ such that $|P_{j}\cap[kn+1]|>k$. Without loss of generality, let us assume $|P_{n}\cap[kn+1]|>k$ and $g_{1},g_{2},\ldots g_{k+1}$ are $k+1$ distinct goods in $P_{n}\cap[kn+1]$ such that $g_{1}\leq g_{2}\leq\ldots\leq g_{k+1}$. For $j\in[k+1]$, let $kn-k+j\in P_{a_{j}}$. Now for all $j\in[k+1]$, swap the goods $kn-k+j$ and $g_{j}$; i.e., iteratively do the following:
  • for all $j\in[k+1]$: $P_{a_{j}}\leftarrow P_{a_{j}}\setminus\{kn-k+j\}\cup\{g_{j}\}$ and $P_{n}\leftarrow P_{n}\setminus\{g_{j}\}\cup\{kn-k+j\}$.
Let $P^{\prime}=(P^{\prime}_{1},P^{\prime}_{2},\ldots,P^{\prime}_{n})$ be the final partition. Since $g_{1},\ldots,g_{k+1}$ are $k+1$ goods in $[kn+1]$ in increasing order of index, $g_{j}\leq kn-k+j$. Thus, $v_{i}(g_{j})\geq v_{i}(kn-k+j)$ and after each of these swaps, the value of $P_{a_{j}}$ cannot decrease. Therefore, for all $j\in[n-1]$, $v_{i}(P^{\prime}_{j})\geq v_{i}(P_{j})\geq\mu^{n}_{i}(M)$ and $(P^{\prime}_{1},\ldots,P^{\prime}_{n-1})$ is a partition of a subset of $M\setminus K$ into $n-1$ bundles of value at least $\mu^{n}_{i}(M)$ for $i$. The lemma follows. ∎
Rule $R^{\alpha}_{k}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered. For $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$, there exists an agent $i$ such that $v_{i}(K)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $K$ to agent $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus K$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Corollary 4 (of Lemma 8).
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, for all $k<|M|/n$ and $\alpha\geq 0$, $R^{\alpha}_{k}(\mathcal{I})$ is an $\alpha$-valid reduction.
Definition 7.
We call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}_{k}$-irreducible, if $R^{\alpha}_{k}(\mathcal{I})$ is not applicable. Moreover, we call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}$-irreducible, if for all $k<|M|/|N|$, $R^{\alpha}_{k}(\mathcal{I})$ is not applicable.
Observation 1.
If $\,\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ is ordered and $R^{\alpha}_{k}$-irreducible for some $k<|M|/n$, then $v_{i}(nk+1)<\alpha\cdot\mu_{i}/(k+1)$ for all agents $i\in N$.
Proof.
By irreducibility of $R^{\alpha}_{k}$ for $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$, we have $v_{i}(K)<\alpha\cdot\mu_{i}$ for all $i\in N$. Since $v_{i}(nk+1)\leq v_{i}(nk)\leq\ldots\leq v_{i}(k(n-1)+1)$, we get $v_{i}(nk+1)\leq v_{i}(K)/(k+1)\leq\alpha\cdot\mu_{i}/(k+1)$. ∎
while for any $k<|M|/n$, $R^{\alpha}_{k}$ is applicable do
       $\mathcal{I}\leftarrow R^{\alpha}_{k}(\mathcal{I})$ for an arbitrary $k$ such that $R^{\alpha}_{k}$ is applicable
end while
return $(N,M,\{v_{i}\}_{i\in N})$
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$.
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Algorithm 3 $\mathtt{R-reduce(\mathcal{I},\alpha)}$:
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$.
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Lemma 9.
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance with additive valuations, and fix an agent $i\in N$. Let $g_{1},g_{2}\in M$ be such that $v_{i}(g_{1})+v_{i}(g_{2})\leq\mu^{n}_{i}(M)$. Then $\mu^{n}_{i}(M)\leq\mu^{n-1}_{i}(M\setminus\{g_{1},g_{2}\}).$
Proof.
To prove the lemma, it suffices to find a partition of (a subset of) $M\setminus\{g_{1},g_{2}\}$ into $n-1$ bundles, such that the value of agent $i$ for each bundle is at least $\mu^{n}_{i}(M)$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2}\}\subseteq P_{1}\cup P_{2}$. If $\{g_{1},g_{2}\}\subseteq P_{1}$, then $(P_{2},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[2]$. We have
$v_{i}(P_{1}\cup P_{2}\setminus\{g_{1},g_{2}\})$ $\geq 2\mu^{n}_{i}(M)-\mu^{n}_{i}(M)\geq\mu^{n}_{i}(M).$
Hence $(P_{1}\cup P_{2}\setminus\{g_{1},g_{2}\},P_{3},\ldots,P_{n})$ is a certificate. ∎
Lemma 10.
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and a fixed agent $i\in N$ with additive valuation, let $g_{1},g_{2},g_{3}\in M$ be such that $g_{2}\neq g_{3}$ and $v_{i}(g_{2})+v_{i}(g_{3})\leq\mu^{n}_{i}(M)$. Then $\mu^{n-2}_{i}(M\setminus\{g_{1},g_{2},g_{3}\})\geq\mu^{n}_{i}(M).$
Proof.
To prove the lemma, it suffices to partition (a subset of) $M\setminus\{g_{1},g_{2},g_{3}\}$ into $n-2$ bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup\ldots\cup P_{j}$ for $j\leq 3$. If $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup P_{2}$, then $(P_{3},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[3]$. We have
$v_{i}(P_{2}\cup P_{3}\setminus\{g_{2},g_{3}\})=v_{i}(P_{2})+v_{i}(P_{3})-(v_{i}(g_{2})+v_{i}(g_{3}))\geq 2\mu^{n}_{i}(M)-\mu^{n}_{i}(M)\geq\mu^{n}_{i}(M).$
Hence $(P_{2}\cup P_{3}\setminus\{g_{2},g_{3}\},P_{4},\ldots,P_{n})$ is a certificate. ∎
Rule $S^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{1}$-irreducible. There exists an agent $i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus\{1,n+1\}$. If there exists agent $j\neq i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$: - Duplicate good $1$ and allocate $\{1,n\}$ to agent $j$. - Set $N^{\prime}\leftarrow N^{\prime}\setminus\{j\}$ and $M^{\prime}\leftarrow M^{\prime}\setminus\{n\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 11.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha\leq 1$, $S^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
Proof.
For all agents $i$ who is assigned a bundle $B$ in $S^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. First consider the case that $S^{\alpha}$ removes two agents $i\neq j$. Since the reduction rule $R^{\alpha}_{1}$ is not applicable, for all the remaining agents $a$, we have $v_{a}(\{n,n+1\})<\alpha\cdot\mu^{n}_{a}(M)\leq\mu^{n}_{a}(M)$. Hence, by lemma 10, $\mu^{n-2}_{a}(M\setminus\{1,n,n+1\})\geq\mu^{n}_{a}(M)$.
Now consider the case that $S^{\alpha}$ removes only one agent $i$. Let $a\neq i$ be a remaining agent after allocating $\{1,n+1\}$ to $i$. We have $v_{a}(\{1,n+1\})<\alpha\cdot\mu^{n}_{i}(M)\leq\mu^{n}_{i}(M)$, otherwise, $S^{\alpha}$ would allocate $\{1,n\}$ to $a$ (or some agent other than $i$). By Lemma 9, $\mu^{n-1}_{a}(M\setminus\{1,n+1\})\geq\mu^{n}_{a}(M)$. ∎
Rule $T^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{k}$-irreducible for $k\leq 1$. There exists three different agents $i,j,\ell$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$, $v_{j}(\{2,n+2\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$, and $v_{\ell}(\{1,n+3\})\geq\alpha\cdot\mu_{\ell}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Allocate $\{2,n+2\}$ to $j$. Duplicate good $1$ and allocate $\{1,n+3\}$ to $\ell$. Set $N^{\prime}\leftarrow N\setminus\{i,j,\ell\}$ and $M^{\prime}\leftarrow M\setminus\{1,2,n+1,n+2,n+3\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. If $\alpha\leq 4/5$, for all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 12.
Let $\mathcal{I}$ be a fair division instance with additive valuations and let $S=\{h_{1},h_{2},s_{1},s_{2},s_{3}\}\subseteq M$ be such that for $j\in[2]$, $v_{i}(h_{j})\leq 4/5\cdot\mu^{n}_{i}(M)$ and for $j\in[3]$, $v_{i}(s_{j})\leq 2/5\cdot\mu^{n}_{i}(M)$ for all $i\in N$. Then for all $i\in N$ we have $\mu^{n-3}_{i}(M\setminus S)\geq\mu^{n}_{i}(M).$
Proof.
To prove the lemma, it suffices to partition (a subset of) $M\setminus S$ into $n-3$ bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $S\subseteq P_{1}\cup\ldots\cup P_{k}$ for $k\leq 5$. If $S\subseteq P_{1}\cup P_{2}\cup P_{3}$, then $(P_{4},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, we consider two cases:
  • Case 1: $S\subseteq P_{1}\cup\ldots\cup P_{4}$. We have
    $v_{i}(P_{1}\cup\ldots\cup P_{4}\setminus S)$ $\geq\mu^{n}_{i}(M)(4-2\cdot 4/5-3\cdot 2/5)>\mu^{n}_{i}(M).$
    Therefore, $(P_{1}\cup\ldots\cup P_{4}\setminus S,P_{5},\ldots,P_{n})$ is a certificate.
  • Case 2: $S\subseteq P_{1}\cup\ldots\cup P_{5}$. Let $P_{j}\cap S=\{s_{j}\}$ for $j\in[3]$ and $P_{j+3}\cap S=\{h_{j}\}$ for $j\in[2]$. We have
    $v_{i}(P_{1}\cup P_{2}\setminus S)\geq\mu^{n}_{i}(M)(2-2\cdot 2/5)>\mu^{n}_{i}(M),$
    and
    $v_{i}(P_{3}\cup P_{4}\cup P_{5}\setminus S)\geq\mu^{n}_{i}(M)(3-2\cdot 4/5-2/5)=\mu^{n}_{i}(M).$
    Thus, $(P_{1}\cup P_{2}\setminus S,P_{3}\cup P_{4}\cup P_{5}\setminus S,P_{6},\ldots,P_{n})$ is a certificate. ∎
Lemma 13.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha\leq 4/5$, $T^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
Proof.
For all agents $i$ who is assigned a bundle $B$ in $T^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Note that by the preconditions of the reduction $T^{\alpha}(\mathcal{I})$, $\mathcal{I}$ is $R^{\alpha}_{k}$-irreducible for $k\leq 1$. Thus, by Observation 1, for all the remaining agents $i$, we have $\alpha\cdot\mu^{n}_{i}(M)\geq v_{i}(1)\geq v_{i}(2)$ and $\alpha/2\cdot\mu^{n}_{i}(M)\geq v_{i}(n+1)\geq v_{i}(n+2)\geq v_{i}(n+3)$. Since $\alpha\leq 4/5$, by Lemma 12, $v^{n-3}_{i}(M\setminus\{1,2,n+1,n+2,n+3\})\geq v^{n}_{i}(M)$. ∎
Remark 1.
In subsequent sections, similar to Definition 7, an ordered instance is $S^{\alpha}$-irreducible and $T^{\alpha}$-irreducible, if $S^{\alpha}$ and $T^{\alpha}$ are not applicable respectively.
5.3 Combining Round Robin and Bag-Filling
In this section, we present Algorithm 4 ($\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$) which combines the idea of bag-filling and round robin to achieve an $\alpha$-MMS allocation with copies for ordered instances.
Note that the input $\mathcal{I}$ of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ is $R^{\alpha}$-irreducible. For all the agents $i$, we define $\mu_{i}=\mu^{n}_{i}(M)$ to be the MMS value of agent $i$ in the beginning of the Algorithm 4. We do a round robin fashion process as following: starting from $n$ empty bags $B_{1},\ldots,B_{n}$, for $j$ in the cyclic sequence $1,2,\ldots,n$, we add the most valuable available good (not yet allocated to any bag) to $B_{j}$. If for some remaining agent $i$, $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$, then we allocate $B_{j}$ to $i$ and remove $i$ from the set of agents and $j$ from the cyclic sequence. We call this phase of the algorithm, the round robin phase. Note that for all the remaining bags $B_{a}$ and $B_{b}$, if $a<b$, then $v_{i}(B_{a})\geq v_{i}(B_{b})$ for all agents $i$ and hence, first index $1$ will be removed, then index $2$ and so on (see Lemma 15). Let $\{B_{a},\ldots,B_{n}\}$ be the set of the remaining bags at the end of the round robin phase. In the next phase of the algorithm, called duplication phase, we duplicate goods $11,2,\ldots,n-a+$ 1, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. Unlike the round robin phase, it is not clear whether the agents who are allocated during the duplication phase value their bundle at least $\alpha$ fraction of their MMS. However, in Sections 5.4 and 5.5, we show that by appropriately reducing the instance in advance, we can ensure both the desired approximation guarantees and the bound on the number of distinct copies.
Let $B_{i}=\emptyset$ for $i\in[|N|]$
$n\leftarrow|N|$, $j\leftarrow 1$, $a\leftarrow 1$
for $g\leftarrow 1$ to $|M|$ do $\triangleright$ Round robin Phase
       $B_{j}\leftarrow B_{j}\cup\{g\}$
       if there exists $i\in N$ such that $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$ then
             $A_{i}\leftarrow B_{j}$ for such an arbitrary $i$
             $N\leftarrow N\setminus\{i\}$, $a\leftarrow a+1$
       end if
      $j\leftarrow j+1$
       if $j>n$ then
             $j\leftarrow a$
       end if
end for
for $j\leftarrow a$ to $n$ do $\triangleright$ Duplication Phase
       Let $i\in N$
       $A_{i}\leftarrow B_{j}\cup\{j-a+1\}$, $N\leftarrow N\setminus\{i\}$
end for
return $(A_{1},A_{2},\ldots,A_{n})$
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An $\alpha$-MMS allocation $A$ with copies
Algorithm 4 $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$:
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An $\alpha$-MMS allocation $A$ with copies
For the rest of this (sub-)section, let $\alpha\geq 0$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$ without loss of generality.
Lemma 14.
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the bags $B_{j}$ with $|B_{j}|\geq 3$ that were allocated to other agents during the round robin phase, we have $v_{i}(B_{j})\leq 4\alpha/3$.
Proof.
Let $g$ be the last good added to $B_{j}$. We have $g\geq 2n+1$ and hence $v_{i}(g)\leq v_{i}(2n+1)\leq\alpha/3$. The last inequality holds since the instance is $R^{\alpha}_{2}$-irreducible (by Observation 1). We have $v_{i}(B_{j}\setminus\{g\})<\alpha$. Otherwise, $B_{j}\setminus\{g\}$ would have been allocated to $i$, or some other agent who values it at least $\alpha$, previously. Therefore we have,
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)\leq\alpha+\alpha/3=4\alpha/3.\qed$
Lemma 15.
Let $\mathcal{B}$ be the set of bags allocated during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$. Then =$\mathcal{B}=\{B_{1},\ldots,B_{|\mathcal{B}|}\}$.
Proof.
We prove that first $B_{1}$ gets allocated, then $B_{2}$ and so on. Towards contradiction, assume for some $j>1$ bag $B_{j}$ gets allocated to agent $i$ while bag $B_{j-1}$ is not assigned. At the time that $B_{j}$ gets assigned, we have $|B_{j-1}|=|B_{j}|$ since both $B_{j-1}$ and $B_{j}$ were present in all the rounds so far. Let $B_{j-1}=\{g_{1},\ldots,g_{k}\}$ and $B_{j}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k}\}$. Followed from round robin, for all $\ell\in[k]$, we have $g_{\ell}=g^{\prime}_{\ell}-1$ which means $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell})$. Hence $v_{i}(B_{j-1})\geq v_{i}(B_{j})\geq 1$ which is a contradiction that $B_{j-1}$ was not assigned to anyone before moving to $B_{j}$. ∎
5.4 $6/7$-MMS with Copies
In this section, we present an algorithm which given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/2\rfloor$ distinct copies for any $\alpha\leq 6/7$.
First we reduce the instance in Algorithm 5 using $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $S^{\alpha}$. It is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of the algorithm, then in total at most $\lfloor k/2\rfloor$ duplicated goods are introduced, each good is duplicated at most once, and all the instances of these goods (original and copy) are allocated to different agents. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of them can be used. Since all these rules are $\alpha$-valid reductions, it suffices to give a simple $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/2\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
Then, we run Algorithm 4 ($\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$) on the reduced instance. We prove that by the end of the round robin phase, at most $\lfloor n/2\rfloor$ agents remain. In other words, if the bags $B_{a},B_{a+1},\ldots,B_{n}$ are the remaining bags, then $a\geq\lfloor n/2\rfloor+1$. Furthermore, we prove for all remaining agents $i$, $v_{i}(B_{j}\cup\{n-a+1\})\geq\mu_{i}(\mathcal{I})$ for all $j\in\{a,\ldots,n\}$. Therefore, to ensure MMS for the remaining agents, it suffices to duplicate goods $1,2,\ldots,n-a+1\leq\lfloor n/2\rfloor<a$, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. This is exactly the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$. In this phase, we duplicate at most $\lfloor n/2\rfloor$ many goods and at most one copy of each good.
$\mathcal{I}\leftarrow\mathtt{R-reduce(\mathcal{I})}$
while $S^{\alpha}(\mathcal{I})$ is applicable do
       $\mathcal{I}\leftarrow S^{\alpha}(\mathcal{I})$
       $\mathcal{I}\leftarrow\mathtt{R-reduce(\mathcal{I})}$
end while
return $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Algorithm 5 $\mathtt{S-reduce(\mathcal{I},\alpha)}$:
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Lemma 16 follows from Lemma 11 and Corollary 4.
Lemma 16.
For an arbitrary ordered instance $\mathcal{I}$ and $\alpha\leq 1$, let $\mathcal{I}^{\prime}=\mathtt{S-reduce(\mathcal{I},\alpha)}$. Then for all agents $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
For the rest of this (sub)-section, let $\alpha\leq 6/7$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $S^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Observation 2.
For all bags $B_{j}\in\mathcal{B}$, $|B_{j}|\geq 3$.
Proof.
Since $\mathcal{I}$ is $S^{\alpha}$-irreducible, $v_{i}(\{j,n+j\})\leq v_{i}(\{1,n+1\})<\alpha$ for all $i\in N$. Thus, if $B_{j}$ was allocated to some agent, $|B_{j}|>2$. ∎
Lemma 17.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>1\}|>n^{\prime}$.
Proof.
Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>1\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq 1$. By Observation 2, for all $B\in\mathcal{B}$ , $|B|\geq 3$. By Lemma 14, for all $B\in\mathcal{B}^{+}$, $v_{i}(B)\leq 4\alpha/3$. Also, let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$). Therefore, we have
$n$ $\leq v_{i}(M)$
$=\sum_{B\in\mathcal{B}^{+}}v_{i}(B)+\sum_{B\in\mathcal{B}\setminus\mathcal{B}^{+}}v_{i}(B)+\sum_{B\in\bar{\mathcal{B}}}v_{i}(B)$
$<|\mathcal{B}^{+}|4\alpha/3+(|\mathcal{B}|-|\mathcal{B}^{+}|)+(n-|\mathcal{B}|)\alpha$
$\leq(|\mathcal{B}|-|\mathcal{B}^{+}|)+6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B^{+}}|)$ ($\alpha\leq 6/7$)
Removing $(|\mathcal{B}|-|\mathcal{B}^{+}|)$ from both sides of the inequality, we obtain,
$n-|\mathcal{B}|+|\mathcal{B}^{+}|$ $<6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B}^{+}|)=2|\mathcal{B}^{+}|/7+6/7(n-|\mathcal{B}|+|\mathcal{B}^{+}|).$
Therefore, $|\mathcal{B}^{+}|>(n-|\mathcal{B}|+|\mathcal{B}^{+}|)/2$ which means $|\mathcal{B}^{+}|>n-|\mathcal{B}|=n^{\prime}$. ∎
Corollary 5 (of Lemma 17).
$n^{\prime}\leq\lfloor n/2\rfloor$.
Lemma 18.
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq 1$.
Proof.
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 17, there exists at least $n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>1$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
$v_{i}(B_{j}\cup\{n^{\prime}\})$ $=v_{i}(B_{j})+v_{i}(n^{\prime})\geq v_{i}(B_{j})+v_{i}(i_{n^{\prime}})=\sum_{\ell\in[k]}v_{i}(g_{\ell})+v_{i}(i_{n^{\prime}})$
$\geq\sum_{\ell\in[k^{\prime}-1]}v_{i}(g^{\prime}_{\ell+1})+v_{i}(i_{n^{\prime}})=v_{i}(B_{i_{n^{\prime}}})\geq 1.\qed$
Theorem 4.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n^{*}$ agents and $\alpha\leq 6/7$,
$\mathsf{BagFill\mbox{-}RoundRobin}(\mathtt{S-reduce(\mathcal{I},\alpha)},\alpha)$
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies.
Proof.
Let $n$ be the number of agents after applying $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Since $S^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{S-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. $g\leq n^{\prime}$ and thus $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 18, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/2\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Only when applying $S^{\alpha}$, new duplications are introduced and one duplication per two agents that are removed. So during $\mathtt{S-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/2\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $S^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/2\rfloor$. Therefore, in total at most $\lfloor n^{*}/2\rfloor$ distinct copies are used and the allocation is simple. ∎
5.5 $4/5$-MMS with Copies
In this section, we present an algorithm which, given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/3\rfloor$ distinct copies for any $\alpha\leq 4/5$.
Similar to Section 5.4, first we reduce the instance but this time by running Algorithm 6 which uses $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $T^{\alpha}$. Analogous to Algorithm 5, it is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of Algorithm 6, then (i) at most $\lfloor k/3\rfloor$ duplicated goods are introduced, (ii) each good is duplicated at most once, and (iii) all the duplicated goods and their corresponding original instances are allocated to distinct agents. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of these goods can be used and no other goods can be allocated to these agents. Since all these rules are $\alpha$-valid reductions, it suffices to give a simple $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/3\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
Then, we run Algorithm 4 ($\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$) on the reduced instance. Following a similar idea, but a more involved calculation, we prove that we end up with a $4/5$-MMS allocation using at most $\lfloor n/3\rfloor$ distinct copies.
$\mathcal{I}\leftarrow\mathtt{R-reduce(\mathcal{I})}$
while $T^{\alpha}(\mathcal{I})$ is applicable do
       $\mathcal{I}\leftarrow T^{\alpha}(\mathcal{I})$
       $\mathcal{I}\leftarrow\mathtt{R-reduce(\mathcal{I})}$
end while
return $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Algorithm 6 $\mathtt{T-reduce(\mathcal{I},\alpha)}$:
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
Output: An ordered instance $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ with $N^{\prime}\subseteq N$ and $M^{\prime}\subseteq M$.
Lemma 19.
For an arbitrary ordered instance $\mathcal{I}$ and $\alpha\leq 4/5$, let $\mathcal{I}^{\prime}=\mathtt{T-reduce(\mathcal{I},\alpha)}$. Then for all agents $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Proof.
The lemma follows from Lemma 13 and Corollary 4. ∎
For the rest of this (sub)-section, let $\alpha\leq 4/5$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $T^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Lemma 20.
For all bags $B_{j}\in\mathcal{B}$ with $j>2$, $|B_{j}|\geq 3$.
Proof.
Towards contradiction, assume for some $j>2$, $|B_{j}|\leq 2$ and $B_{j}$ is allocated to agent $i$. Note that since $\mathcal{I}$ is $R_{0}^{\alpha}$-irreducible, $v_{a}(1)<\alpha$ for all agents $a$ and hence $|B|\geq 2$ for all $B\in\mathcal{B}$. So for all $\ell\in[n]$, we have $\{\ell,n+\ell\}\in B_{\ell}$ and $B_{j}=\{j,n+j\}$. Note that for all $j^{\prime}<j$, $v_{i}(\{j^{\prime},n+j^{\prime}\})>v_{i}(\{j,n+j\})\geq\alpha$. Hence, for all $j^{\prime}<j$, $|B_{j^{\prime}}|=2$. Let $a_{1},a_{2},a_{3}$ be the three agents who received $B_{1},B_{2},B_{3}$ respectively. Note that the precondition of Rule $T^{\alpha}$ is met for these agents which is a contradiction with $\mathcal{I}$ being $T^{\alpha}$-irreducible. ∎
Observation 3.
For $n\geq 2$, $\{B_{1},B_{2}\}\subseteq\mathcal{B}$ and $v_{i}(B_{j})\leq 3\alpha/2$ for all $i\in N$ and $j\in[2]$.
Proof.
Since the instance is $R^{\alpha}_{0}$-irreducible, $v_{i}(2)\leq v_{i}(1)<\alpha$ for all $i\in N$. Therefore, $|B_{j}|\geq 2$. Let $g$ be the last good added to $B_{j}$. Since $g\geq n+1$, $v_{i}(g)\leq v_{i}(n+1)<\alpha/2$. We have
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)<\alpha+\alpha/2=3\alpha/2.$
Now assume $B_{2}\notin\mathcal{B}$. Let $i$ be an agent who is not allocated any bundle in $\mathcal{B}$. Note that for all $B\notin\mathcal{B}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated during the round robin phase. Therefore
$n$ $\leq v_{i}(M)$
$=\sum_{j}v_{i}(B_{j})$
$<3\alpha/2+(n-1)\alpha$
$\leq(4/5n+2/5)\cdot,$ ($\alpha\leq 4/5$)
which is a contradiction with $n\geq 2$. ∎
Lemma 21.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}|\geq 2n/3$ or $n\leq 5$.
Proof.
Assume $n\geq 6$. Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\alpha$. Let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ many remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). We consider two cases:
Case 1: $v_{i}(B_{2})\leq 4\alpha/3$.
For all $B\in\mathcal{B}\setminus\{B_{1},B_{2}\}$, Lemma 20 implies $|B|\geq 3$ and Lemma 14 implies $v_{i}(B)\leq 4\alpha/3$. Also by Observation 3, $v_{i}(B_{1})\leq 3\alpha/2$ and we have $v_{i}(B_{2})\leq 4\alpha/3$. Therefore
$\sum_{B\in\mathcal{B}^{+}}v_{i}(B)$ $\leq 4\alpha/3(|\mathcal{B}^{+}|-1)+3\alpha/2$
$\leq(16|\mathcal{B}^{+}|/15+2/15).$ ($\alpha\leq 4/5$)
We have
$n$ $\leq v_{i}(M)$
$=\sum_{B\in\mathcal{B}^{+}}v_{i}(B)+\sum_{B\in\mathcal{B}\setminus\mathcal{B}^{+}}v_{i}(B)+\sum_{B\in\bar{\mathcal{B}}}v_{i}(B)$
$\leq(16|\mathcal{B}^{+}|/15+2/15)+(|\mathcal{B}|-|\mathcal{B}^{+}|)\alpha+(n-|\mathcal{B}|)\alpha$
$\leq(2/15+4n/5+4|\mathcal{B}^{+}|/15).$ ($\alpha\leq 4/5$)
We obtain
$n/5$ $\leq 4|\mathcal{B}^{+}|/15+2/15.$
Therefore, $|\mathcal{B}^{+}|\geq(3n-2)/4\geq 2n/3$ (for $n\geq 6$).
Case 2: $v_{i}(B_{2})>4\alpha/3$.
If $|B_{2}|\geq 3$, then by Lemma 14, $v_{i}(B_{2})\leq 4\alpha/3$. Also if $|B_{2}|=1$, $v_{i}(B_{2})<\alpha$ since the instance is $R^{\alpha}_{0}$ irreducible. Therefore, $|B_{2}|=2$. It implies $|B_{1}|=2$ and we have $B_{1}=\{1,n+1\}$ and $B_{2}=\{2,n+2\}$. Let $a$ and $b$ be the agents who received $B_{1}$ and $B_{2}$. If $v_{i}(\{1,n+3\})\geq\alpha$, then the instance is not $T^{\alpha}$-irreducible since $a,b,i$ satisfy its precondition. Therefore, $v_{i}(1)+v_{i}(n+3)<\alpha$.
$v_{i}(1)+v_{i}(n+2)$ $\geq v_{i}(2)+v_{i}(n+2)=v_{i}(B_{2})>4\alpha/3.$
Therefore,
$v_{i}(n+3)$ $<v_{i}(n+2)-\alpha/3$
$<(\alpha/2-\alpha/3)$ ($v_{i}(n+2)\leq v_{i}(n+1)<\alpha/2$ due to Observation 1)
$=\alpha/6.$
Now fix an index $j>2$ and let $g$ be the last good added to $B_{j}$. We have
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)$
$<\alpha+\alpha/6$
$=7\alpha/6.$
Therefore,
$n$ $\leq v_{i}(M)$
$=v_{i}(B_{1})+v_{i}(B_{2})+\sum_{B\in\mathcal{B}\setminus\{B_{1},B_{2}\}}v_{i}(B)+\sum_{B\in\bar{\mathcal{B}}}v_{i}(B)$
$<3\alpha+(|\mathcal{B}|-2)7\alpha/6+(n-|\mathcal{B}|)\alpha$
$\leq(4n/5+(2|\mathcal{B}|+8)/15)$ ($\alpha\leq 4/5$)
$\leq(14n/15+4/15).$ ($|\mathcal{B}|\leq n-2$)
We obtain that $n<4$ which is a contradiction with $n\geq 6$. ∎
Corollary 6 (of Lemma 21).
$n^{\prime}\leq\lfloor n/3\rfloor$ for $n\geq 6$.
Lemma 22.
If $n\geq 6$, for all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq\alpha$.
Proof.
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 21, there exists at least $2n/3\geq n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\alpha$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
$v_{i}(B_{j}\cup\{n^{\prime}\})$ $=v_{i}(B_{j})+v_{i}(n^{\prime})$
$\geq v_{i}(B_{j})+v_{i}(i_{n^{\prime}})$
$=\sum_{\ell\in[k]}v_{i}(g_{\ell})+v_{i}(i_{n^{\prime}})$
$\geq\sum_{\ell\in[k^{\prime}-1]}v_{i}(g^{\prime}_{\ell+1})+v_{i}(i_{n^{\prime}})$
$=v_{i}(B_{i_{n^{\prime}}})>\alpha.$
Theorem 5.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n^{*}>5$ agents and $\alpha\leq 4/5$,
$\mathsf{BagFill\mbox{-}RoundRobin}(\mathtt{T-reduce(\mathcal{I},\alpha)},\alpha)$
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies.
Proof.
Let $n$ be the number of agents after applying $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Since $T^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{T-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their value. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. Since $g\leq n^{\prime}$, $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 22, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now give an upper bound on the total number of copies and prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/3\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Only when applying $T^{\alpha}$, new duplications are introduced and one duplication per three agents that are removed. So during $\mathtt{T-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/3\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $T^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/3\rfloor$. Therefore, in total at most $\lfloor n^{*}/3\rfloor$ distinct copies are used and the allocation is simple. ∎
Appendix A Full MMS by Removing $n-2$ Bads for Chore Allocation
Consider the case where for an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $i\in N$ and $j\in M$, $v_{ij}$ is the cost of performing chore $j$. We assume agents are additive, that is the cost of bundle $S\subseteq M$ for agent $i\in N$ equals $v_{i}(S)=\sum_{j\in S}v_{ij}.$ Agents want to minimize their cost.
Recall that we denote the collection of all allocations of the set $M$ among $n$ agents by $\mathcal{A}_{n}(M)$. The maximin share [^4] Although the maximin share is defined with a min-max type formula, if one considers chores to be goods with negative value, then a max-min formulation is equivalent. (MMS) of agent $i$ is the upper bound on the cost agent $i$ may exert by splitting the set $M$ into $n$ bundles, getting the worst out of them. Formally:
Definition 8 (Maximin share (MMS)).
For an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n$ agents, the maximin share (MMS) of agent $i$ with cost function $v_{i}$, denoted by $\mu_{i}^{n}(M)$, is given by
$\mu_{i}^{n}(M)=\min_{A\in\mathcal{A}_{n}(M)}\max_{S\in A}v_{i}(S).$
For every agent $i$, we normalize $\mu_{i}^{n}(M)=1$, thus $v_{i}(M)\leq n$ for every $i\in N$. We next present Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}$, which operates in an analogous way to Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill}$ from Section 4.1.
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be the MMS partition of agent 1
Consider the bipartite graph $G=(V=N\times P,E)$, where $E=\{(i,P_{j})\ :\ v_{i}(P_{j})\leq 1\}$
if there is a perfect matching between $N$ and $P$ in $G$ then
       Allocate each $i\in N$ its match in the perfect matching
end if
else
       Let $P^{\prime}\subset P$ be the minimal set of bundles violating Hall’s theorem ($|P^{\prime}|>1$ since $1$ is matched to every node in $P$)
       Consider some $P_{j}\in P^{\prime}$ and let $\hat{P}=P^{\prime}\setminus\{P_{j}\}$
       Find a matching between bundles in $\hat{P}$ and agents in $N$
       foreach $i\in N$ matched to some $P_{j}\in\hat{P}$ do
             $A_{i}\leftarrow P_{j}$
       end foreach
      Let $\tilde{N}$ be the set of unallocated agents and $\tilde{M}=M\setminus\left(\bigcup_{i\in N\setminus\tilde{N}}A_{i}\right)$ the set of unallocated items
       Allocate agents in $\tilde{N}$ using procedure $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}((\tilde{N},\tilde{M},\{v_{i}\}_{i\in\tilde{N}}))$
end if
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $m-n+2$ allocated items.
Algorithm 7 $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $m-n+2$ allocated items.
$t\leftarrow 0$, $B_{t}\leftarrow\emptyset$
$N_{0}\leftarrow N$, $M_{0}\leftarrow M$
for items $j\leftarrow 1$ to $m$ do
       if $\exists i\in N_{t}$ such that $v_{i}(B_{t}\cup\{j\})<1$ then
             $B_{t}\leftarrow B_{t}\cup\{j\}$
       end if
      else
             Let $i_{t}\in\operatorname*{arg\,min}_{i\in N_{t}}v_{i}(B_{t})$
             $A_{i_{t}}\leftarrow B_{t}$
             $t\leftarrow t+1$, $B_{t}\leftarrow\emptyset$
             $N_{t}\leftarrow N_{t-1}\setminus\{i_{t-1}\}$, $M_{t}\leftarrow M_{t-1}\setminus(B_{t-1}\cup\{j\})$
       end if
end for
Let $i_{t}$ be the last remaining agent in $N_{t}$. Set $A_{i_{t}}\leftarrow B_{t}$.
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $|M|-|N|+1$ allocated items.
Algorithm 8 $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}(\mathcal{I}=(N,M,\{v_{i}\}_{i\in N}))$:
Input: $n$ additive cost functions $\{v_{i}\}_{i\in N}$ over items in $M$.
Output: Allocation $A=(A_{1},\ldots,A_{n})$ with at least $|M|-|N|+1$ allocated items.
We first show that $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}$ never gives an agent more than their MMS value, and throws away at most $|N|-1$ items.
Lemma 23.
Assuming that for every $i\in N$, $v_{i}(M)\leq|N|$, $\mathsf{BagFill\mbox{-}and\mbox{-}Remove}$ outputs and allocation such that $v_{i}(A_{i})\leq 1$ for every $i\in N$, and allocates at least $|M|-|N|+1$ items.
Proof.
Every time an agent is allocated, except the last agent, we throw away the current item considered. Thus, we allocate all items but $|N|-1$. Since every time we allocate, we remove a set of items that is worth at least 1 for all agents from $M_{t}$, we maintain the invariant that $v_{i}(M_{t})\leq|N|-t$. Thus, when we reach the last agent, they get a bundle of cost at most 1. ∎
We now show that at most $n-2$ items should be discarded.
Lemma 24.
For any fair division instance of chores $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with additive costs, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}(\mathcal{I})$ outputs a full MMS allocation and allocates at least $|M|-|N|+2$ items.
Proof.
As in the proof of Lemma 4, if all agents are allocated by a perfect matching in $G$, and then we do not need to throw away any items. If we match a subset of the agents by bundles in $G$, then the remaining agents do not have any edges to the allocated bundles, which means that if $k$ agents were allocated, the rest of the agents have a cost of at most $|N|-k$ for the remaining bundles. Thus, we can apply Lemma 23 and get that at most $|N|-k-1$ items are thrown away. Since $k\geq 1$, we get that the number of allocated items is at least $|M|-|N|+2$. ∎
Appendix B Missing Proofs of Section 3
Proof of Lemma 2.
For any agent $i$, let $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$ be his MMS partition. Consider the MMS allocations $A=(A_{1},\dots,A_{n})$ where $A_{i}\sim U(\{P^{i}_{1},\dots,P^{i}_{n}\})$. Namely, $i$’s bundle, $A_{i}$, is one of his MMS bundles chosen uniformly at random, independent of other choices. We use the balls-and-bins analogy [38], and consider the process of constructing $A$ as throwing $n$ balls (the agents) into $n$ bins (the index in an MMS partition). We are interested in the number of balls per bin, so we ignore the fact that the balls are distinct.
Fix some good $g\in M$ and assume without loss that $g\in P^{i}_{1}$ for any $i\in N$. The event in which $g$ is copied at least $k$ times for the allocation $A=(A_{1},\dots,A_{n})$, is exactly the event in which the bin corresponding to index $1$ has at least $k$ balls. A simple analysis shows that,
$\Pr[g\text{ was copied at least }k\text{ times}]\leq\binom{n}{k}\left(\frac{1}{n}\right)^{k}\leq\left(\frac{e}{k}\right)^{k},$
where the last transition follows from Stirling’s approximation.
If we plug in $k=\frac{3\ln m}{\ln\ln m}$, we get
$\left(\frac{e}{k}\right)^{k}$ $=\left(\frac{e\ln\ln m}{3\ln m}\right)^{\frac{3\ln m}{\ln\ln m}}$
$\leq\exp\left(\frac{3\ln m}{\ln\ln m}\ln\left(\frac{e\ln\ln m}{3\ln m}\right)\right)$
$\leq\exp\left(\frac{3\ln m}{\ln\ln m}\left(\ln\ln\ln m-\ln\ln m\right)\right)$
$=\exp\left(-3\ln m+\frac{3\ln m\ln\ln\ln m}{\ln\ln m}\right).$
For large enough $m$ we have $\left(\frac{e}{k}\right)^{k}<\exp(-2\ln m)=\frac{1}{m^{2}}$.
Applying the union bound we have,
Pr[there exists a good with more than $\Pr\left[\text{there exists a good with more than }\frac{3\ln m}{\ln\ln m}\text{ copies}\right]<m\cdot\frac{1}{m^{2}}=\frac{1}{m},$
which completes the proof. ∎
Appendix C Missing Proofs of Section 4
Proof of Lemma 3.
Every time an agent is allocated, except the last agent, we duplicate the last item that was put in his bag and put the duplicate as the first item in the next bag. Since every item’s value is less than $1$ for all agents, every allocated bundle contains at least two items. Therefore, we duplicate a different item each time, and in total, $|N|-1$ items. We now show that upon termination, every agent gets at least their MMS value.
We show that for every $t$, for every $i\in N_{t}$, $v_{i}(M_{t})\geq|N|-t$. We prove by induction on $t$. Obviously, it’s true for $t=0$. Assume this is true for $t-1\in[|N|-1]$, and consider the iteration when agent $i_{t-1}$ gets assigned bundle $A_{i_{t-1}}$. Consider some agent $i\in N_{t}$. By the inductive assumption, we have that for every $i^{\prime}\in N_{t-1}$, $v_{i^{\prime}}(M_{t-1})\geq|N|-t+1.$ As $N_{t}\subset N_{t-1}$, we have that
$v_{i}(M_{t-1})\geq|N|-t+1.$ (1)
Now consider the bundle $B_{t-1}$ just before some item $j$ was added to it, causing agent $i_{t-1}$’s value to rise above 1. Since before adding $j$, the value each agent in $N_{t-1}$ assigned to $B_{t-1}$ was smaller than 1, we also have
$v_{i}(B_{t-1})<1.$ (2)
This implies that
$v_{i}(M_{t})\ =\ v_{i}(M_{t-1}\setminus B_{t-1})\ =\ v_{i}(M_{t-1})-v_{i}(B_{t-1})>|N|-t,$
where the last inequality follows Equations (1), (2). ∎
Appendix D MMS with copies for $k$-demand valuations
It is natural to consider other classes of valuation functions, not only arbitrary monotone and additive. One of the simplest valuations is that of 1-demand, every agent only wants one good, and every good has a value. Trivially, one round of round-robin will result in an MMS allocation. So, what about $k$-demand? The valuation of a set is the sum of the $k$ most valuable goods in this set.
It is natural to ask if the algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ is relevant, unfortunately, this does not work. Consider the following input: All $n$ agents with identical valuation functions $v_{i}=v$, $v$ is 2-demand: $v(g)\geq 0$ for all goods, and $v(S)=\max_{g,h\in S}v(g)+v(h)$ for $|S|\geq 2$.
We now construct an example on which algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ fails, we define goods and valuations as follows: there are $n-1$ goods $g_{1},\ldots,g_{n-1}$ for which $v(g_{j})=(1-\epsilon)/2$, $n-1$ goods $h_{1},\ldots,h_{n-1}$ for which $v(h_{j})=(1+\epsilon)/2$, and two additional goods $x_{1}$, $x_{2}$ for which $v(x_{j})=1/2$. One possible MMS allocation is $\{g_{1},h_{1}\}\ldots\{g_{n-1},h_{n-1}\},\{x_{1},x_{2}\}$, and the MMS value is one.
Now consider [arbitrarily] adding goods in the order $g_{1},\ldots,g_{n-1}$ to a bag $B$, at this point in time $v(B)<1$. Next, add $x_{1}$ and $x_{2}$ to $B$, all agents have value 1 for the bag, if the bag is assigned to one of the agents not enough goods will remain. Duplicating the good $x_{2}$ and all of the $h_{j}$ will not suffice to get an MMS allocation for the remaining $n-1$ agents.
This suggests a variant of algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ for $k$-demand valuations. Upon processing a new good $g$, and adding it to $B$, do not assign the entire bag $B$ to an agent with $v(B)\geq 1$. Instead, choose some set $S\subseteq B$, $|S|=k$ with $v(S)\geq 1$. Some such set $S$ must exist (as $v(B)\geq 1$ and $v$ is $k$-demand). Moreover, the last good added to the bag $g$ must be in $S$, otherwise the bag $B^{\prime}=B\setminus\{g\}\subset B$, considered just prior to the arrival of $g$, would also have had value $\geq 1$. Give one of the agents with $v(B)\geq 1$ some $k$-subset of $B$, $S\subset B$, for which $v(S)\geq 1$. Return all goods in $B\setminus S$ to the pool of goods yet to be processed, and assign $B=\{g\}$, the last good added to that triggered this allocation, this creates a copy of $g$. Also, such duplications can be done at most $n-1$ times and no good is duplicated more than once.
To Summarize:
Observation 4.
For $k$-demand valuations $n-1$ distinct copies suffice to achieve an MMS allocation.
Appendix E Example for Algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$
Agent 1 Agent 2 Agent 3 Agent 4
$g_{1}$ 0.2 0.3 0.8 0.1
$g_{2}$ 0.2 0.5 0.3 0.1
$g_{3}$ 0.7 0.3 0.2 0.8
$g_{4}$ 0.3 0.2 0.3 0.2
$g_{5}$ 0.3 0.8 0.1 0.2
$g_{6}$ 0.5 0.5 0.1 0.7
$g_{7}$ 0.1 0.1 0.5 0.1
$g_{8}$ 0.1 0.1 0.7 0.2
$g_{9}$ 0.8 0.3 0.3 0.8
$g_{10}$ 0.2 0.8 0.1 0.3
$g_{11}$ 0.2 0.2 0.9 0.3
$g_{12}$ 0.7 0.3 0.1 0.5
Figure 2: Running example for $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$. There is an MMS allocation without copies for this instance. All MMS values are one. For agent 1 we have $v_{1}(\{g_{1},g_{2},g_{3}\})=1.1$, $v_{1}(\{g_{4},g_{5},g_{6}\})=1.1$, $v_{1}(\{g_{7},g_{8},g_{9}\})=1$ and $V_{1}(\{g_{10},g_{11},g_{12}\})=1.1$. For agent 2 we have $v_{2}(\{g_{1},g_{2},g_{3}\})=1.1$, $v_{2}(\{g_{4},g_{5}\})=1$, $v_{2}(\{g_{6}.g_{7},g_{8},g_{9}\})=1$ and $V_{2}(\{g_{10},g_{11},g_{12}\})=1.1$. By construction the MMS partitions consist of contiguous runs of $g_{1},\ldots,g_{12}$.
Consider the case when we add items in order $g_{4},g_{7},g_{9},g_{1},g_{2},g_{10},g_{12},g_{7},g_{5},g_{3},g_{8}\ldots$
  • For $B=\{g_{4},g_{7}\}$, $v_{1}(B)=0.4$, $v_{2}(B)=0.3$, $v_{3}(B)=0.8$, and $v_{4}(B)=0.3$, no agent has attained the MMS value for this set $B$.
  • For $B=\{g_{4},g_{7},g_{9}\}$, $v_{1}(B)=1.2$, $v_{2}(B)=0.8$, $v_{3}(B)=1.1$, and $v_{4}(B)=1.1$, arbitrarily choose agent 1 and assign $B$ to agent 1, set $B$ to be a copy of the last good added: $B=\{g_{9}\}$.
  • For $B=\{g_{9},g_{1},g_{2}\}$, we have $v_{2}(B)=1.1$, $v_{3}(B)=0.8$, $v_{4}(B)=1$, arbitrarily choose agent 2 and assign $B$ to agent 2, set $B=\{g_{2}\}$.
  • For $B=\{g_{2},g_{10},g_{12},g_{7}\}$, we have $v_{3}(B)=1.2$, $v_{4}(B)=1$ and assign $B$ to agent 4, set $B=g_{7}$.
  • For $B=\{g_{7},g_{5},g_{3},g_{8}\}$ we have that $v_{3}(B)=1.5$ and assign $B$ to agent 3.
Note that not all goods have been allocated, we could add the remaining goods to any of the agents, arbitrarily.
The execution of $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ depends on the order in which the items are added.
Appendix F Picking Sequence Fails for Allocation with Copies
Consider three agent $\{a,b,c\}$ with additive valuations over three goods $x,y,z$. Let the ordinal valuations be as follows:
$v_{a}:$ $x\succ z\succ y$
$v_{b}:$ $x\succ z\succ y$
$v_{c}:$ $x\succ y\succ z$
Consider the ordered counterpart (see Definition 4) of this instance and the allocation $A^{\prime}$ below, in which every good is copied once.
$A^{\prime}_{a}=\{1,2\},\qquad A^{\prime}_{b}=\{1,3\},\qquad A^{\prime}_{c}=\{2,3\}$
We interpret this (virtual) allocation as inducing the following value guarantee: for each good $j$ that agent $i$ receives in $A^{\prime}$, she is entitled to a good whose value is at least her $j$ most-valuable good in the original instance. E.g., agent $b$ in the above must receive $x$ and either of $y$ or $z$.
One can generalize the picking sequence algorithm ([16], see also Proposition 2), to a setting with copies in two ways. Assume we have a virtual allocation $A^{\prime}$, where each good has at most $k$ extra copies.
One way is to run $k+1$ instances of the picking sequence algorithm, where in each run each good in $M$ is allocated at most once. The first agent to pick is an agent which holds (one of the instances of) the virtual good $1$, the second holds $2$ and so on. Another way, is to allow all agents which hold good $1$ to pick first, then all agents which hold good $2$, and so on. In this interpretation agent $i$ picks the most-valued good $g$ such that $g\notin A_{i}$, and $g$ was allocated no more than $k+1$ times (so each original good is copied at most $k$ times as well). Below we show that neither of these interpretations work if we break the ties arbitrarily. In fact, the first one fails for any tie-breaking rule, see Example 2.
The failure of the two generalizations follows from the fact that an agent is not allowed to have two instances of the same good. If valuations were such that any $q$ copies of any good $g$ would yield a value of $q\cdot v(g)$, the same proof as in [16] would carry over to this setting. We show examples in which the two generalizations fail when at most one extra copy is made (i.e., for $k=1$). For the first generalization, the example below shows that breaking ties arbitrarily may fail.
Example 1.
Consider the valuations and (virtual) allocation presented at the beginning of the section. Assume ties are broken such that the picking sequences in the two rounds are $a,c,b$ and $b,a,c$, respectively. The first run results in $(A_{a},A_{b},A_{c})=(\{x\},\{z\},\{y\})$. In the second run, since the picking sequence is $b,a,c$, before agent $c$ gets to pick the allocation would be $(A_{a},A_{b},A_{c})=(\{x,z\},\{z,x\},\{y\})$ which leaves no item for $c$ to pick in the last round.
For the second generalization, consider the following example, in which the value guarantees do not hold, regardless of tie-breaking.
Example 2.
$v_{a}:$ $x\succ y\succ z$
$v_{b}:$ $z\succ y\succ x$
$v_{c}:$ $z\succ y\succ x$
with the allocation
$A^{\prime}_{a}=\{1,2,3\},\qquad A^{\prime}_{b}=\{1\},\qquad A^{\prime}_{c}=\{2,3\}$
Agents which hold good 1 pick: $(A_{a},A_{b},A_{c})=(\{x\},\{z\},\emptyset)$.
Agents which hold 2 pick: $(A_{a},A_{b},A_{c})=(\{x,y\},\{z\},\{z\})$.
Agents which hold 3 pick: Agent a cannot pick another good.